### v1

# Extremely luminous optical afterglow of a distant and energetic gamma-ray burst GRB 230204B

Abstract
Robotic telescope networks play an important role in capturing early and bright optical afterglows, providing critical insights into the energetics and emission mechanisms of GRBs. In this study, we analyze GRB 230204B, an exceptionally energetic and multi-pulsed long GRB, detected by the Fermi GBM and MAXI detectors, with an isotropic equivalent gamma-ray energy exceeding 1054 erg. Time-resolved spectral analysis reveals a transition in the prompt emission from hard (sub-photospheric dominated) spectra during early pulses to softer (synchrotron radiation dominated) spectra in later pulses, indicative of a hybrid jet composition. We report the discovery and characterization of the optical afterglow using the MASTER and BOOTES robotic telescope networks, alongside long-term radio observations extending to 335 days post-burst with the ATCA. At 1.3 ks post-burst, the optical luminosity was exceptionally high, surpassing even other bright GRBs, such as GRB 221009A (the “BOAT”). Multi-wavelength modeling, incorporating data from MASTER, BOOTES, DOT, Swift/XRT, and radio observations, was conducted using an external ISM forward-shock top-hat jet model with afterglowpy. The results reveal a narrow and highly collimated jet with a circumburst density of 28.12 cm-3, kinetic energy 4.18 erg, and a relatively low value of = 2.14 , indicating shock-compression of magnetic field in the surrounding interstellar medium. We constrained a low radiative efficiency of 4.3 %. This study highlights the indispensable contribution of robotic networks to early afterglow observations and advances our understanding of GRB 230204B unique characteristics and underlying jet physics.
1 Introduction
Gamma-ray bursts (GRBs) are among the most energetic events in the Universe, emitting vast amounts of energy within a few seconds to minutes, often followed by an afterglow that can be observed across the electromagnetic spectrum (Piran, 1999; Kumar & Zhang, 2015). The discovery of afterglows in the late 1990s marked a significant milestone in GRB research, enabling precise localization and redshift determination, which, in turn, provided insights into the distances and energies involved (Costa et al., 1997; van Paradijs et al., 1997; Frail et al., 1997). The study of GRB afterglows, particularly in the optical regime, has been pivotal in understanding the progenitor systems, the environment surrounding the burst, and the mechanisms driving the emission (Kann et al., 2010; Li et al., 2012; Gupta, 2023). The temporal evolution of these afterglows, often characterized by a power-law decay, can reveal crucial details about the jet dynamics, energy injection, and interaction with the surrounding medium (Meszaros & Rees, 1997a; Sari et al., 1998).
In recent times, robotic telescope networks, such as Mobile Astronomical System of Telescope-Robots (MASTER), Burst Observer and Optical Transient Exploring System (BOOTES), and others, have become indispensable in the rapid follow-up of GRBs, often identifying optical afterglows within minutes to hours (especially for Fermi Gamma-ray Burst Monitor (GBM) detected GRBs with large error circles) of the initial burst (Lipunov et al., 2010a; Castro-Tirado et al., 2012). Fermi GBM’s localization accuracy, while adequate for high-energy studies, often leaves large localization uncertainties spanning a few to tens of degrees radius, making rapid and wide-field follow-up observations essential. These robotic telescope networks, equipped with wide-field cameras and fully automated systems, are specifically designed to address this challenge. Their ability to quickly respond to GRB alerts and scan large areas of the sky allows them to pinpoint transient optical emissions within the Fermi error regions. Additionally, their prompt response times enable them to observe early afterglow phases, providing data at these early times critical for informing the onset and evolution of the emission (Gupta et al., 2021b; Ror et al., 2023b; Sánchez-Ramírez et al., 2024). While these observations are useful for constraining GRB redshifts indirectly by identifying afterglow candidates for follow-up spectroscopic studies, they typically rely on single-band detections and are limited in sensitivity, restricting their ability to detect only the bright afterglows. Despite these limitations, the robotic nature of these observatories ensures continuous and autonomous temporal coverage, improving the chances of detecting GRBs in real-time and providing crucial data to guide further multi-wavelength follow-up efforts (Castro-Tirado, 2023).
GRB 230204B, detected by the Fermi GBM and MAXI, stands out as one of the most energetic GRBs observed, with an isotropic equivalent energy (Eγ,iso) exceeding 1054 erg. The optical afterglow of this burst was discovered by the MASTER and BOOTES robotic telescope networks approximately 1.3 ks post-burst, marking the beginning of the broadband follow-up campaign. Additionally, at the time of MASTER/BOOTES detection ( 1.3 ks post-burst), the optical luminosity was found to be exceptionally high when compared to other bright bursts, suggesting that GRB 230204B harbored an extremely luminous afterglow phase. In this paper, we present a detailed analysis of the optical and multi-wavelength follow-up observations (see Section 2) of GRB 230204B, including data from MASTER, BOOTES, Devasthal Optical Telescope (DOT), Swift Ultraviolet and Optical telescope (UVOT) and radio telescope. We explore the temporal and spectral characteristics of the prompt emission and afterglow and compare its optical luminosity with other known GRBs, the results are presented in Section 3. Discussion on potential physical mechanisms driving this energetic burst and the afterglow brightness comparison are given in Section 4. The key findings are summarized in Section 5.
2 Observations and data analysis
This section details the observation campaign for GRB 230204B, observed by multiple space-based and ground-based telescopes across a broad wavelength range, from gamma rays to radio frequencies. Comprehensive details of the observing campaign are provided below and the sequence of prompt and afterglow observations for GRB 230204B is shown in Figure 1.
2.1 Prompt Emission
GRB 230204B was first detected on 4th February 2023 by the Fermi satellite at 21:44:27.20 UT (hereafter T0), with duration of 216 seconds in 50-300 keV (Poolakkil et al., 2023). MAXI also detected it at 21:47:51 UT and the burst position was constrained to approximately RA = 13h 10m 19s, Dec = -21° 45′ 07′′ (J2000) with a statistical uncertainty region of 0.1°(Serino et al., 2023). The burst did not trigger the Neil Gehrels Swift Observatory; henceforth Swift (Gehrels et al., 2004). However, following the MAXI trigger notification, the Gamma-ray Urgent Archiver for Novel Opportunities (GUANO, Tohuvavohu et al. 2020) pipeline requested data from the onboard buffer of the Swift Burst Alert Telescope (BAT) and initiated a search within the 200-second event mode data centered around T0. The GRB was detected in this Swift/BAT GUANO data, determining its to be approximately 65 seconds (Kennea et al., 2023). The burst was also detected by the Indian multiwavelength satellite AstroSat, which measured a of seconds (Waratkar et al., 2023). The prompt emission of the burst was also seen by AGILE (Casentini et al., 2023) and GRBAlpha (Dafcikova et al., 2023) missions, with every mission providing a slightly different value of duration as expected due to the different sensitive and spectral coverage.
2.2 Prompt X-ray Emission: MAXI
MAXI (mounted on the ISS) scans a position of the sky every 92 minutes (Matsuoka et al., 2009). The position of the GRB moved into the MAXI field of view at T0 +192.8 sec, and the scan transit ended at T0 +248.8 sec. During the scan transit, the MAXI/GSC nova alert system (Negoro et al., 2016) triggered the burst at 21:47:51 UT (T0 + 203.8 sec), with MAXI’s detection revealing the soft nature of the burst. Assuming the source flux was constant over the transit, the X-ray flux in the 4.0–10.0 keV band was calculated to be 628 49 mCrab (1 sigma error, Serino et al. 2023).
After the detection of prompt emission, MAXI/GSC could not observe the position of the burst for about 6 hours due to the observation constraint of the radiation zone (Sugizaki et al., 2011) in the International Space Station (ISS) orbit. The observation resumed at T0 + 22.5 ks, and there were 11 consecutive scan transits before the start of the observations by Swift/XRT. The afterglow was not detected in any transit. We calculated 3 sigma upper limits in the 2–20 keV band (see Table 8 in appendix) for these transits following the method by Sugita et al. (2018).
2.3 Afterglow: Soft X-ray Emission
The fading X-ray counterpart of GRB 230204B was detected by Swift/XRT approximately 80 ks after the burst trigger (D’Elia et al., 2023). Given the faint nature of the X-ray afterglow at such a late epoch, all observations were performed in photon counting (PC) mode to optimize sensitivity for faint sources. Follow-up observations were conducted during two epochs (XRT exposure times of 5.1 ks, and 4.5 ks, respectively), on 5 February 2023 and 14 February 2023, resulting in only two usable data points (see Table A2 of the appendix). The final observation on February 14 approached the detection limit of the instrument, emphasizing the faintness of the X-ray emission and the challenges of constraining the X-ray afterglow’s properties at these late times.
2.4 Afterglow: Optical Emission
We discovered the optical afterglow of GRB 230204B using two robotic telescope networks (using MASTER in the unfiltered band and using BOOTES in the clear band, respectively) independently (see Figure 1). The optical counterpart of the burst was also identified by the 0.7m GROWTH-India Telescope (GIT) approximately 1.57 hours (later than MASTER and BOOTES) after the burst trigger (Swain et al., 2023). This optical afterglow was subsequently observed by several other ground-based telescopes (Smartt et al., 2023; Ror et al., 2023a). The redshift of the burst, = 2.142, was determined through spectroscopic observations using the X-shooter spectrograph on the Very Large Telescope (VLT, Saccardi et al. 2023). The given magnitude (for the positive detection) and upper limits within 3 are listed in Table LABEL:tab:optical.
2.4.1 MASTER robotic telescope
MASTER global network consists of 9 identical fully robotic wide-field twin telescopes with 4-8 square degrees field of view, distributed around the Earth for all-sky monitoring up to 19-20 mag with the goal of detecting optical transients, including high energy astrophysics events such as GRBs, Gravitational waves, ultra-high energy neutrino and Fast Radio Bursts (Lipunov et al., 2010b, 2019, 2022, 2023). We observed such events as soon as possible upon receiving the alerts (Ershova et al., 2020) using the GCN system (Barthelmy et al., 1998; Barthelmy, 2003, 2016) every night in both, the fully robotic open (8 square degrees FOV) and closed (4 square degrees FOV) modes, being ready for observations with fully clear hemisphere (clamshell dome constructions). Open mode is used in the case of alerts provided by Fermi, LIGO/Virgo, IceCube, ANTARES/KM3Net (with large error boxes) when at least 2 images per night are required if large error boxes are at least 5 degrees above the horizon, in order to cover as much as possible with our given field of view. Closed mode (using both tubes in parallel Kornilov et al. (2012)) is preferred for events with error boxes smaller than 1 degree such as the ones provided by Swift, MAXI, Fermi-LAT, GECAM, Einstein Probe, INTEGRAL, when polarization observations can be undertaken (Troja et al., 2017; Jordana-Mitjans et al., 2020; Gupta et al., 2022a). The first exposure duration also depends on the time between the trigger and the start of observation and can be increased from 10 sec at the first minute to 180 sec at later stages in order to detect the fainter objects.
GRB 230204B was observed by MASTER-Kislovodsk and MASTER-SAAO. First images were received for the MAXI error box, then telescopes automatically pointed to Fermi field. The MASTER-Kislovodsk robotic telescope started observing the MAXI trigger 979708795 observation at 2023-02-04 22:05:55 UT (1064 sec after trigger time which was 74 sec after the notification time, see Figure 1) at 11∘ altitude, in 44∘ distance to the Moon (0.3 phase, that creates significant interference for wide-field instruments observations) through clouds at horizon with =14.8 mag at first image in polarization filter.
The MASTER-SAAO robotic telescope located in the South African Astronomical Observatory, started observing error box automatically 85 seconds after notice time (i.e. 1075 sec after trigger time) at 2023-02-04 22:06:06 UT with optical counterpart MASTER OT J131034.94-214304.8 detection, coincident with the counterpart reported by GIT, ATLAS, VLT, DOT, and BOOTES111https://gcn.gsfc.nasa.gov/other/230204B.gcn3. The observations began at 27∘ altitude with an exposure time of 180 sec in the polarization filter. The OT unfiltered magnitude in the first image was 12.9 mag (=17.6 mag). We used a reference image on 2020-04-24.82 UT with unfiltered =20.6 mag.
For the photometric analysis of this object, catalog data for comparison stars were sourced from Gaia EDR3 to evaluate the variance in magnitude errors. The photometry dataset excludes frames where the comparison star data were deemed unreliable. The magnitude error for comparison stars was computed as the deviation between the weighted average magnitude (where weights are the inverse square of the signal-to-noise ratio) and the magnitude measured on individual frames. Two approaches were employed to determine the stellar magnitude. In the first approach, a reference star from Gaia EDR3 was identified, its flux on the observational frame measured, and the magnitude calculated using Pogson’s scale. In the second approach, a zero-point correction was applied to the reference star’s magnitude, utilizing the comparison stars as secondary references. The second method generally produced a smaller magnitude variance compared to the first, owing to the improved calibration achieved through the zero-point correction. This dual-approach methodology ensures robust photometric accuracy while mitigating systematic errors.
2.4.2 BOOTES robotic telescope
BOOTES followed up GRB 230204B with three of the six 60 cm robotic telescopes, namely the ones at the BOOTES-4/MET station in Lijiang Astronomical Observatory (Yunnan, China), at BOOTES-5/JGT station in Observatorio Astronómico Nacional of San Pedro Mártir (Baja California, México) and at the BOOTES-2/TELMA station in La Mayora (Málaga, Spain). Following the MAXI trigger, the BOOTES-4/MET telescope performed two epoch observations at 2023-02-04 22:19:12 UT and 2023-02-05 17:42:04 UT. The BOOTES-2/TELMA telescope and the BOOTES-5/JGT telescope also executed follow-up observations at 2023-02-05 04:16:55 UT and 2023-02-05 07:43:45 UT, respectively. A series of images were obtained using the clear filter with exposures of 1 sec, 5 sec, and 10 sec at BOOTES-4/MET telescope during the first epoch and the optical afterglow was clearly detected during the decaying phase. However, during the following multi-filters (Sloan- and clear filters, 60 sec exposure) monitoring, the optical afterglow was not detected on the stacked images for each epoch (as they were not as deep as the clear filter exposures). All the images were processed with bias, dark subtraction, and flat-field correction using custom IRAF routines (Gupta et al., 2021b). Photometry was carried out using the standard IRAF package and the images were calibrated with nearby comparison stars from the USNO-B1.0 and SDSS catalogs.
2.4.3 DOT
We initiated observations using the DOT (Pandey, 2016; Gupta et al., 2024a) on 2023-02-06 at 01:49:47.939 UT, approximately 1.16 days after the trigger. Multiple frames with an exposure time of 180 sec each were observed in the , , and bands, and continuous observations were conducted over four consecutive nights using the same filters. For image preprocessing, we employed the standard IRAF framework, and DAOPHOT II was used to determine magnitudes from the cleaned files. We stacked the images for each night in the same filter to increase the signal-to-noise ratio. PSF photometry was performed on the final stacked images (Gupta et al., 2022b, c). Due to the high moon phase during the burst observation, we only detected GRB 230204B in the and -band images from the first night. For the remaining filters, we could only establish upper limits. Photometric calibration is performed using the standard stars from the Pan-STARRS catalog.
2.4.4 Swift/UVOT
The Swift/UVOT began settled observations of the field of GRB 230204B 80.6 ks after the MAXI trigger (Serino et al., 2023). No optical afterglow was detected by UVOT. Photometry was obtained from co-added UVOT image mode data. The source counts were extracted from the UVOT stacked images using a source region of 5 arcsec radius. Background counts were extracted using an aperture of 20 arcsec in radius located in a source-free region near the GRB. We use the Swift tool uvotsource to obtain the background subtracted count rates using source and background apertures of 5 and 20 arcseconds in radius, respectively. They were converted to magnitude using the UVOT photometric zero points (Breeveld et al., 2011; Poole et al., 2008).
2.4.5 Radio Emission
We observed GRB 230204B on 8 February 2023 (3.6 days post-burst), with the Australia Telescope Compact Array (ATCA, Project code CX527) at a wide range of frequencies centered on 5.5, 9.0, 16.7, 21.2, 33 and 35 GHz, each with a 2048 MHz-wide band. We reduced the visibility data using standard routines in MIRIAD (Sault et al., 1995). We used a combination of manual and automatic RFI flagging before calibration, conducted with MIRIAD tasks uvflag and pgflag, respectively. We used PKS 193463 to determine the bandpass response for frequency bands 5.5 and 9 GHz, and PKS 0727115 as the bandpass calibrator for 16.7, 21.2, 33 and 35 GHz bands. We used PKS 193463 to calibrate the flux density scale for all frequency bands. We used PKS 1256220 to calibrate the time-variable complex gains for all epochs and frequency bands. After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (Clark, 1980) on the target source field using standard MIRIAD tasks INVERT, CLEAN and RESTOR to obtain the final images.
We detect the radio counterpart at 16.7 and 21.2 GHz at a position consistent with the BOOTES/MASTER/GIT optical counterpart position (Swain et al., 2023). For all other frequencies, we report 5-sigma upper limits (see Table 7 of appendix). The GRB position was also observed with the Australian SKA Pathfinder by the Rapid ASKAP Continuum Survey (RACS; McConnell et al., 2020) on 2023 July 1 and 2024 Jan 5. We find no radio detections in preliminary data and report 5-sigma upper limits of 1.5 mJy. There were also no archival radio detection pre-burst within 1 arcminute in surveys: the National Radio Astronomy Observatory VLA Sky Survey (Condon et al., 1998), the Sydney University Molonglo Sky Survey (Mauch et al., 2003), the Very Large Array Sky Survey (Lacy et al., 2020) or the RACS survey. We measured a VLASS 5-sigma upper limit of 0.69 mJy at 3 GHz.
2.4.6 Swift BAT Survey Data
Swift/XRT did not observe the location of X-ray afterglow of GRB 230204B until 91.5 ks post-burst. Therefore to search for the hard X-ray early afterglow, we used the BatAnalysis python package (Parsotan et al., 2023) to analyze Swift BAT survey data from 2023-02-03 to 2023-02-11. We analyzed survey data, where the location of the GRB had at least a partial coding fraction of on the BAT detector plane. As expected, the GRB afterglow was not detected in the BAT survey data, which is typical since hard X-ray afterglows have never been detected at such late times. However, we were able to place 5 upper limits on the emission in the 14-195 keV energy band (shown in Figure 2). We additionally conducted a mosaiced analysis of the BAT survey data in 1-day time bins, for two days after the trigger time, and were able to place 5 flux upper limits on the emission in the 14-195 keV energy band where the flux of the GRB afterglow is erg/sec/cm2.
Based on the above observations, we have shown the broadband composite afterglow light curve of GRB 230204B in Figure 2.
3 Results
3.1 Prompt emission temporal and spectral characterize
For the analysis of Fermi GBM data, we utilized the Fermi GBM Data Tools222https://fermi.gsfc.nasa.gov/ssc/data/analysis/gbm/gbm_data_tools/gdt-docs/ to extract temporal information. We selected specific NaI (NaI 7 and NaI 8) and BGO (BGO 1) detectors for the analysis based on their optimal viewing angles relative to the GRB’s position to maximize the signal-to-noise ratio. For the NaI detectors, we focused on two energy bands: 8-30 keV and 50-300 keV, allowing us to explore both low and high-energy photon populations. For BGO 1, two bands were used: 0.3-1 MeV and 1-40 MeV, to investigate higher-energy photons. The prompt emission background-subtracted light curve of GRB 230204B, obtained using Fermi GBM across different energy channels, reveals a complex multi-pulse structure, as shown in Figure 3. This GRB exhibits four distinct emission episodes, each separated by periods of quiescence, with the shaded regions highlighting the intervals selected for time-integrated spectral analysis. The energy-dependent light curves show that the temporal structure and intensity vary significantly across different energy bands, reinforcing the diverse nature of GRB light curves. Notably, the hardness ratio (HR) demonstrates evolution over time, indicating spectral changes during the burst. The inset plot shows the MAXI light curve, covering the 2-20 keV energy range, further subdivided into 2-4 keV, 4-10 keV, and 10-20 keV. Although MAXI detected only part of the GRB, its observations provide valuable insights into the lower-energy emission and suggest a very softer nature of the last pulse of GRB 230204B.
We analyzed the time-integrated (T0-0.80 to T0 +229.69 sec) Fermi spectrum of GRB 230204B using the Multi-Mission Maximum Likelihood framework (3ML) to evaluate its spectral characteristics. We explore several phenomenological models, including the Band function, Cutoff power-law (CPL), and combinations with a Blackbody component, which is theoretically expected to originate from the photospheric emission of GRBs. For each model, we employed Bayesian spectral fitting to obtain the Deviance Information Criterion (DIC) values, allowing us to systematically compare the model fits. The results indicate that the Band function provides a statistically better fit to the time-integrated spectrum compared to the CPL alone, as indicated by the DIC values. Furthermore, the addition of a Blackbody component alongside the Band function yielded an even lower DIC value, suggesting that this model combination better describes the observed spectral features.
In addition to the time-integrated spectrum, we analyzed the spectral evolution across four distinct emission episodes of GRB 230204B, fitting each episode with the same set of phenomenological models. Our findings show that the first and third episodes are best described by the CPL + Blackbody model, while the second and fourth episodes favor the Band + Blackbody combination based on DIC values. The DIC values and the best-fit spectral parameters for both the time-integrated analysis and individual episodes are summarized in Table 1 and Table 2. These results highlight the diversity in spectral shapes across the burst duration and provide valuable insights into the complex emission mechanisms underlying GRB 230204B, supporting the presence of a thermal component in combination with non-thermal models.
| Time-intervals | DIC value | |||
| (from T0 in sec) | Band | CPL | BB+Band | BB+CPL |
| -0.80-229.69 | 7118.17 | 7123.59 | 7094.00 | 7103.16 |
| -0.80-4.97 | 3245.26 | 3245.68 | 3220.72 | 3214.46 |
| 37.00-68.15 | 5083.02 | 5091.67 | 5058.11 | 5077.71 |
| 109.450-161.90 | 5655.64 | 5660.37 | 5619.57 | 5606.26 |
| 172.58-229.69 | 5601.18 | 5606.86 | 5575.45 | 5577.01 |
| Time-intervals | Best-fit parameters | |||||
| (from T0 in sec) | (keV) | (keV) | kT (keV) | |||
| -0.80-229.69 | -0.78 | 519.44 | -3.37 | – | – | 15.08 |
| -0.80-4.97 | – | – | – | -0.38 | 259.10 | 24.98 |
| 37.00-68.15 | -0.65 | 661.98 | -2.61 | – | – | 15.51 |
| 109.450-161.90 | – | – | – | -0.70 | 514.31 | 14.01 |
| 172.58-229.69 | – | – | – | -0.80 | 239.09 | 10.62 |
GRBs are known for their significant spectral evolution and high variability, with emission properties often varying notably across individual pulses within multi-pulse bursts. Li et al. (2021) conduct a comprehensive time-resolved Bayesian spectral analysis of well-defined GRB pulses observed by Fermi. They observe a trend toward spectral softening over time, largely reflected in the gradual decrease of the low-energy power-law index. We compared the time-resolved spectral analysis of GRB 230204B with the sample studied by Li et al. (2021). We also observe that the low-energy power-law indices of GRB 230204B become softer over time (hard to soft), with a distinct spectrum evolution from the first to later episodes (see Figure 4). Such spectral evolution feature of prompt emission could be explained as photospheric emission is most commonly observed near the trigger time, while synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Li et al., 2021).
3.2 Pulse-wise prompt emission correlation and analysis
The pulse-wise analysis of GRB prompt emission provides a valuable perspective on the Amati and Yonetoku correlations (key relationships used to explore the physical properties of GRBs), enhancing our understanding of GRB spectral evolution and energetics on a finer timescale (Chand et al., 2020). By examining each pulse independently, rather than averaging over the entire prompt emission phase, we gain insights into the distinct characteristics of individual emission episodes within a single burst. In the Amati correlation, which relates the isotropic energy () to the spectral peak energy in the rest frame (Amati, 2006), each pulse may fall along or diverge from the established long and short GRB trends, providing information about the pulse-specific energy release mechanisms. Such an analysis not only clarifies the evolutionary trajectory of the GRB emission but also tests the applicability of the Amati correlation at sub-burst levels (Basak & Rao, 2013).
In the top-left panel of Figure 10, we illustrate the positioning of each pulse (plotted as colored squares) of GRB 230204B within the established Amati correlation. For reference, samples of long and short bursts from Minaev & Pozanenko (2020) are represented by orange and blue circles, respectively. The linear fits corresponding to these groups are indicated by solid orange and blue lines, with 3 uncertainties depicted by shaded bands around each fit. The time-integrated and pulse-wise analysis of GRB 230204B follows the Amati plane of long GRB populations. The top-right panel of Figure 10 presents GRB 230204B positions within the Yonetoku correlation framework, a relation between peak energy and the isotropic peak luminosity () in the rest frame (Yonetoku et al., 2004). Here, the short and long GRB samples from Nava et al. (2012) are shown by blue and orange circles, with the 3 scatter around each population’s best-fit line visualized by shaded areas. GRB 230204B follows the Yonetoku correlation and lies on the upper side of this correlation. The results in both correlations offer an insightful examination of GRB 230204B, highlighting the pulse-specific properties within established GRB behavior and indicating the energetic and luminous nature of the burst.
3.2.1 Spectral Peak-Duration Distribution Analysis
Long-duration GRBs are generally characterized by a softer spectrum relative to the typically harder spectrum observed in the majority of short-duration GRBs, as demonstrated in previous studies (e.g., Kouveliotou et al. 1993). To analyze this characteristic for GRB 230204B, we measured both the time-integrated and episode-wise spectral peak energy (an indicator of spectral hardness) alongside the duration (the time during which 90% of the burst’s total fluence is detected, for example, calculation for the second pulse of GRB 230204B is shown in the bottom left of Figure A1 of the appendix). We then plotted the pulse-wise values on the - plane to compare with a larger dataset of GRBs detected by the Fermi GBM, as illustrated in the bottom right panel of Figure 10.
Our results indicate that the time-integrated and episode-wise properties of GRB 230204B exhibit softer values and durations exceeding 2 seconds, placing it on the right side of the spectral peak-energy versus duration plane. This position aligns well with the properties typically observed for long-duration GRBs, as they occupy the region in the - plane corresponding to longer durations and softer spectra. This comparison confirms that GRB 230204B follows the standard characteristics of long-duration GRBs.
3.3 Afterglow analysis and correlation
Based on the multi-wavelength follow-up observations campaign ( 1.3 ks to 335 days post burst) of GRB 230204B, we noted that BOOTES robotic and DOT observations have a well-sampled light curve in optical / filters. We fitted flux density optical /-band light curves using a simple power-law function with a temporal decay index of -1.85 0.03 (/DOF = 43.47/52). Furthermore, we combined all the optical data (MASTER, BOOTES, DOT, and a few points from GCN) to get a more well-sampled light curve and it could also be fitted with a simple power-law function with a temporal decay index of -1.84 0.02 (/DOF = 81.13/58). We noted that the optical temporal index is steeper than the typically observed decay slope of -1 (Dainotti et al., 2022). On the other hand, the X-ray afterglow light curve of GRB 230204B has very limited (close to the sensitivity of XRT) data points as Swift followed the burst during the late phase and could be fitted with a temporal decay index of -0.72 0.63 (/DOF = 0.58/1).
The correlation between the average optical decay index and optical luminosity provides crucial insights into the physical mechanisms governing GRB afterglows and their surrounding environments. Oates et al. (2012, 2015) examined the possible correlation between the average optical temporal decay indices and optical luminosity (at 200-sec post burst) for GRBs with optical afterglows observed by Swift/UVOT. They found evidence of anti-correlation between these two parameters (200-sec post burst) as shown in Figure 5. This correlation indicates that optical luminous bursts decay more rapidly than optical faint bursts. We measured optical luminosity and temporal decay index for GRB 230204B and overlaid it onto the luminosity-decay correlation. GRB 230204B exhibits a relatively steep decay rate ( 1.4 in MASTER filter) and high optical luminosity (at the first epoch of afterglow detection). We noted GRB 230204B, highlighted in red in the figure, aligns well with the global correlation observed for long-duration GRBs and lies towards the edge (see Figure 5). Possible physical explanations of this correlation are discussed in section 4.2.
3.4 Broadband afterglow modeling of GRB 230204B
| Model | Top hat | Gaussian | Gaussian Core |
| 0.003 | 0.004 | 0.005 | |
| 0.006 | 0.006 | 0.007 | |
| – | 4.020 | 4.714 | |
| 55.621 | 55.529 | 55.143 | |
| 2.280 | 2.267 | 2.252 | |
| 1.449 | 1.987 | 2.685 | |
| -5.669 | -5.899 | -6.133 | |
| -0.525 | -0.353 | -0.198 | |
| 1.0 | 1.0 | 1.0 | |
| ln () | -101.310.08 | -107.500.08 | -110.530.08 |
| ln (BF) | 7638.16 | 7631.97 | 7628.94 |
The broadband afterglow modeling involved fitting the observed data across multiple wavelengths, including radio, optical, and X-ray bands. Such broadband modeling helps to understand the physical parameters of the jet, such as its opening angle, energy, and the surrounding medium’s density. The afterglow emission is generally explained by the synchrotron radiation model (Sari et al., 1998; Meszaros & Rees, 1997b; Sari & Piran, 1999). In this model, relativistic electrons are accelerated by the GRB shockwave spiral in the magnetic field, emitting radiation across the electromagnetic spectrum. The synchrotron spectrum is characterized by several break frequencies (self-absorption, peak, and cooling frequencies), which evolve with time and affect the observed light curves at different wavelengths.
For the afterglow modeling of GRB 230204B, we utilized afterglowpy333It currently does not account for certain features, including reverse shock emission, an external wind medium (n r-2), or synchrotron self-absorption. (Ryan et al., 2020)/Redback (Sarin et al., 2024) to explore various jet structure models, including the Tophat, Gaussian, and GaussianCore models. The Bayesian inference framework provided by Redback allowed us to test different priors and achieve robust parameter estimation. Specifically, we employed the PyMultiNest sampler to explore the parameter space and used well-defined priors for key parameters such as the observer angle (), core angle (), initial energy (E0), and interstellar medium density (nISM). Due to the limited number of data points, we have fixed bulk Lorentz factor ( 800), constrained using the -Lorentz factor correlation (Liang et al., 2010). The priors used in the analysis include uniform distributions for the energy, core angle, ambient medium density, and a sine prior for the observer’s viewing angle.
The Top-Hat model provided the best fit to the data (see Figure 6), with log evidence (ln ) of -101.31 0.08 and a Bayes factor (ln BF) of 7638.16 as indicated by the Bayesian evidence comparison. These results strongly suggest that the jet structure for GRB 230204B is best described by a simple Top-Hat profile (see Table 3). The posterior distributions (see corner plot as shown in Figure 11 of the appendix) constrained viewing angle of 0.003 rad and a core angle 0.006 rad (limit as no jet break is observed), suggesting a narrow, highly collimated jet structure. The inferred initial energy is relatively high at 55.62 erg, and the ISM density is constrained to 1.45 cm-3.
The Top-Hat jet model provides a robust description of GRB 230204B’s afterglow, indicating a narrow, highly collimated jet. This is consistent with the high fluence observed in the early afterglow phase. The tightly constrained parameters for and indicate that we are likely observing the jet from a viewing angle close to the core, providing a relatively straightforward interpretation in the context of synchrotron emission from a top-hat jet model. A notable feature in the fit is the very low value of , the magnetic field fraction, which was constrained to = -5.67. This low magnetic field can plausibly arise from shock compression of the ambient interstellar medium (ISM), which supports a weakly magnetized medium (Kumar & Barniol Duran, 2009; Wang et al., 2015). Similar values of have been observed in GRBs such as GRB 190530A (Gupta et al., 2022a), GRB 171205A (Leung et al., 2021), GRB 130427A (Fraija et al., 2016), and a few others (Kumar & Barniol Duran, 2009, 2010; Santana et al., 2014), highlighting the diversity in GRB environments. For GRBs, when , there might be Synchrotron Self-Compton (SSC) contribution to afterglow, particularly at higher frequencies (Joshi & Razzaque, 2021). However, in the case of GRB 230204B, the SSC effects appear negligible within the observed wavelength range (i.e., low-frequencies radio and optical), consistent with findings for similar GRBs (Gupta et al., 2022a; Wang et al., 2024). This is further supported by the temporal decay slopes, afterglow modeling (afterglowpy primarily models synchrotron emission processes), and lack of GeV emission (consistent with synchrotron-dominated regimes). Future studies focusing on a broader sample of bright GRBs with low , combined with well-sampled broadband light curves (from gamma-ray to radio) and afterglow modeling that includes SSC contributions, have the potential to shed light on their physical origins and diverse emission mechanisms.
4 Discussion
4.1 Spectral evolution and Possible Radiation Mechanism
The spectral evolution of the prompt emission reveals complex temporal and spectral variability, which provides crucial insights into the nature of the emission mechanisms and the physical conditions in the GRB jet (Bošnjak & Daigne, 2014). Typically, GRB prompt spectra evolve over time, with the peak energy often shifting to lower energies as the burst progresses. This trend, known as “hard-to-soft” evolution, is commonly observed and is thought to reflect cooling processes within the jet (Norris et al., 1986; Bhat et al., 1994). Alternatively, “tracking” behavior, where follows the intensity variations of the burst, has also been observed in some cases (Ryde & Svensson, 1999), suggesting a possible relationship between the central engine activity and the observed emission (Gupta et al., 2021b).
The radiation mechanisms driving GRBs are fundamentally linked to the spectral index observed during the prompt emission phase and the overall radiative efficiency. The spectral index , which describes the shape of the GRB’s spectrum, provides crucial insights into the underlying physical processes. A harder low energy spectral index (i.e., closer to or exceeding -2/3) typically suggests a dominant photospheric or thermal component, where radiation emerges directly from the photosphere of the relativistic outflow. In contrast, a softer a (less than -2/3) often points to non-thermal processes, such as synchrotron emission, where electrons are accelerated in magnetic fields, producing a broad power-law spectrum (Chattopadhyay et al., 2022; Caballero-García et al., 2023).
To investigate the spectral evolution of GRB 230204B, we conducted a time-resolved spectral analysis, selecting time bins based on a Bayesian algorithm for optimal selection. The top panel of Figure 7 presents the deviation information criterion (DIC) values, comparing the Band function with the CPL for each time-resolved bin. The DIC comparison indicates that the majority of bins—excluding three—are better described by the traditional Band model. In the bottom panel of Figure 7, we illustrate the temporal evolution of key spectral parameters, including , and , as derived from the Band and CPL models for GRB 230204B (see Table LABEL:TRS_Table). These parameters are overlaid on the count-rate light curve of the burst’s prompt emission, revealing that the observed exhibits an intensity-tracking behavior. Harder spectral indices during the first and second episodes might suggest a prominent photospheric component, which often produces harder spectra due to thermal-like emission (Ryde et al., 2010). The softer spectral indices during the later bins of the second, third, and fourth episodes might suggest a synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Li et al., 2021; Acuner & Ryde, 2018).
Additionally, we explored correlations among the spectral parameters, examining (a) log(Flux) versus log(), (b) log(Flux) versus , and (c) log() versus . Pearson correlation analysis was used to quantify the correlation strengths (Pearson correlation coefficient, (r) and to assess the statistical significance (p-value) for each parameter pair. We noted a strong correlation for log(Flux) versus log() of Band and CPL models with correlation coefficients equal to 0.69 and 0.73, respectively. We also noted a moderate correlation for log(Flux) versus of Band model with a correlation coefficient equal to 0.46, but no correlation for log() versus .
The radiation efficiency of GRBs is another key factor, reflecting the fraction of the jet’s kinetic energy converted into gamma rays. It varies significantly depending on the dominant radiation mechanisms within the jet (Wang et al., 2015). In the photospheric model, radiation is produced as photons escape from the optically thick inner regions of the relativistic outflow. Energy is released at the photosphere where the flow becomes transparent, leading to thermal emission. Photospheric models tend to predict higher radiation efficiencies, sometimes exceeding 50-60%, especially if energy dissipation occurs near the photosphere (Lazzati et al., 2009; Beloborodov, 2010; Pe’er & Ryde, 2011; Lazzati et al., 2013). In contrast, the synchrotron models (arising from accelerated particles in internal shocks or magnetic reconnection sites) often predict lower efficiencies, typically in the range of a few percent to–30%, in such cases much of the energy remains in the kinetic form rather than being radiated away (Kobayashi et al., 1997; Spada et al., 2000; Beniamini & Piran, 2013; Kumar & Zhang, 2015). The hybrid model combines thermal photospheric emission with non-thermal synchrotron radiation from shocks or magnetic reconnections within the jet. This model can accommodate both thermal and non-thermal features in GRB spectra. Hybrid models can achieve efficiencies ranging from a few percent to 50% or higher, as they leverage both thermal and non-thermal emission processes (Zhang & Yan, 2011; Zhang & Pe’er, 2009; Vurm et al., 2011). These discussions suggest that GRBs may involve a complex interplay of thermal and non-thermal processes, with the relative contributions of these mechanisms varying across different bursts and even between different phases of the same burst.
The prompt emission and afterglow observations of GRB 230204B present intriguing insights into its jet structure and radiation mechanisms. Our analysis of the prompt emission reveals a spectral hardening during the initial pulses, with a subsequent softening in later pulses, consistent with a hybrid composition (baryonic + Poynting-flux) of Band + Blackbody components. This spectral evolution suggests that the jet may consist of both photospheric (thermal) and synchrotron (non-thermal) components, with the harder spectrum at the start dominated by a photospheric component and later phases increasingly influenced by synchrotron radiation (Sharma et al., 2020; Gupta et al., 2024c). Additionally, our afterglow modeling indicates a low gamma-ray efficiency of 4.3%, suggesting that most of the jet’s energy remains in kinetic form with minimal conversion into prompt gamma-ray emission. While the low gamma-ray efficiency supports a hybrid model, where both thermal photospheric and non-thermal synchrotron components coexist, it does not entirely rule out a synchrotron-only scenario (Rossi et al., 2002). Such a model, without any photospheric thermal emission, could also account for the low efficiency, provided the energy dissipation mechanism predominantly drives non-thermal radiation. However, the observed spectral evolution, with the initial hard component transitioning to softer emission, is more naturally explained by the presence of a photospheric contribution early on. Furthermore, the low gamma-ray efficiency aligns well with the characteristics of an ISM-like top-hat jet structure in the afterglow phase. In such a scenario, a top-hat jet interacting with a homogeneous interstellar medium produces a relatively uniform energy distribution across the jet, reducing the likelihood of highly efficient gamma-ray emission (Sari et al., 1998; Granot & Sari, 2002; Zhang & Kobayashi, 2005). This jet structure also favors an afterglow powered predominantly by forward shock emission, which is consistent with the observed light curve and its steady decay pattern.
The kinetic energy and of GRB 230204B place it within the population of GRBs characterized by relatively low radiation efficiency. The kinetic energy for GRB 230204B, inferred from afterglow modeling, was found to be significantly higher than the isotropic gamma-ray energy, emphasizing that a substantial fraction of the jet’s energy remained in kinetic form and was not radiated as gamma rays. We noted that GRB 230204B has a higher and values than expected from magnetar central engine (2 1052 erg, Metzger et al. 2011). This supports the rapidly rotating black hole as a possible central engine for this energetic and bright burst (Li et al., 2018; Ror et al., 2024). The distribution of radiation efficiencies across the GRB population, as shown in Figure 8, highlights GRB 230204B as the upper edge with a very low efficiency.
4.2 Brightness Comparison to GRB sample and physical scenarios for GRB 230204B
The optical luminosity of GRB 230204B (R/r-band and g-band data) has been compared with the population of GRBs with measured redshifts (Dainotti et al., 2024), providing valuable insights into its optical emission characteristics. The temporal evolution of GRB 230204B’s optical luminosity (see Figure 9) reveals an exceptionally bright initial emission, followed by a decay phase consistent with a forward shock model in a homogeneous interstellar medium (ISM). This places it among the brightest optical GRBs ever observed at the epoch of its detection, making it a standout event in the context of optical afterglow studies. The extreme optical brightness is an order of magnitude higher than other highly energetic bursts ( 1.3 ks post-burst) such as GRB 080319B (Racusin et al., 2008; Pandey et al., 2009), GRB 990123 (Castro-Tirado et al., 1999), GRB 190114C (Misra et al., 2021; Gupta et al., 2021a), and GRB 221009A. Furthermore, we have also examined the optical luminosity of GRB 230204B at the epoch of its detection as a function of redshift for a population of GRBs with measured redshifts (Dainotti et al., 2024). We noted that GRB 230204B has the highest optical luminosity ( 1.1-2.4 ks post-burst) with respect to all known GRBs with measured redshifts at the same epoch (see Figure 9). This suggests that GRB 230204B could represent a highly efficient energy transfer to the optical afterglow, or it might be related to the properties of its surrounding environment, jet structure, or progenitor system, contributing to its extraordinary luminosity.
Additionally, we noted that GRB 230204B follows an anti-correlation between the optical decay index and optical luminosity (Oates et al., 2012, 2015). There are several scenarios where such a correlation could be explained. Continued energy injection from a central engine, such as a magnetar or a rapidly spinning black hole, could initially produce a very bright afterglow (Zhang & Mészáros, 2001). However, if the injection is short-lived or if the engine rapidly loses power, the afterglow could exhibit a steep decay as the additional energy source wanes (Zhang & Mészáros, 2001; Dai & Lu, 1998). Our afterglow analysis did not show evidence of central engine activity (flare, plateau, variability, etc). A reverse shock could also produce a very luminous optical flash (Zhang & Kobayashi, 2005). If the reverse shock is strong, it can dominate the early afterglow. Once the reverse shock fades, the light curve might steepen significantly, leading to a rapid decay. For GRB 230204B, however, our afterglow modeling suggests that a forward shock model could explain the early optical emission. Another possibility involves a structured jet, where a bright core with steep wings might cause a steep decay if the emission transitions from being core-dominated to wing-dominated, or if there is a slight change in the viewing angle (Rossi et al., 2002; Granot & Kumar, 2003). However, our afterglow modeling suggests that a top-hat jet could explain the afterglow of GRB 230204B, so we discarded the structured jet for this burst. If the GRB is in a wind-like medium (density r-2), the early afterglow can be very bright (Chevalier & Li, 2000). A change in the density profile (e.g., transition to a uniform medium) could cause the afterglow to decay steeply (Panaitescu & Kumar, 2002; Starling et al., 2008). However, for GRB 230204B, our afterglow modeling suggests that an ISM-like medium could explain the broadband emission. A jet break occurs when the relativistic beaming angle of the jet becomes wider than the observer’s line of sight (Racusin et al., 2009). This typically results in a steepening of the afterglow light curve (Liang et al., 2008). An achromatic temporal slope of -2 at late times could indicate that the break occurred, however, for GRB 230204B, synchronous broadband data is not available to confirm such behavior, also such an early jet break is not expected so we discarded the possibility. GRB afterglows depend strongly on the observer’s viewing angle relative to the jet’s core (Granot et al., 1999; van Eerten & MacFadyen, 2013). When an observer views the GRB slightly off-axis but still close to the jet core, the observed afterglow appears very bright initially. This is due to the deceleration of the relativistic jet as it interacts with the surrounding medium, causing the beamed emission to broaden and spread into the observer’s line of sight. As the jet slows further and its emission becomes more isotropic, the light curve steepens, showing a rapid decline (Ryan et al., 2015). Such behavior suggests that GRB 230204B is observed close to the jet axis, as indicated by its narrow inferred viewing angle from afterglow modeling. The steep early optical decay and extreme luminosity support this interpretation.
In Figure 12, we present a comparative analysis of the X-ray luminosity light curves of GRB 230204B with a sample of other GRBs that have measured redshifts (Gupta et al., 2024b). The X-ray flux data for these GRBs were sourced from the Swift XRT light curve repository444https://www.swift.ac.uk. To enable a direct comparison, the light curves are presented in the rest frame in the k-corrected co-moving bandpass, allowing for a consistent examination of intrinsic luminosity evolution across the sample. GRB 230204B is characterized by limited data, with only two available data points recorded at a late epoch, depicted in red on the plot. Despite the scarcity of data, the observed luminosity for GRB 230204B during this late phase is notably consistent with the average luminosity values observed in the comparative GRB sample at a similar epoch. This suggests that GRB 230204B’s X-ray luminosity profile falls within the expected range for GRBs in this phase.
In our analysis, we also examined the radio luminosity of GRB 230204B in the context of a broader sample of GRBs with known redshifts, as compiled by Shilling et al. (in preparation). Figure 12 illustrates the rest-frame 8.5 GHz luminosity light curves for this GRB sample, restricted to measurements with a signal-to-noise ratio (SNR) greater than 2 for clarity. For GRB 230204B specifically, we used a 9 GHz observation from ATCA as a close approximation to the 8.5 GHz rest-frame frequency. This 9 GHz upper limit, taken several days post-burst, is plotted alongside the 8.5 GHz light curve distribution. We noted that GRB 230204B ’s radio upper limit appears on the higher side of the observed luminosity range for GRBs in the sample. The positioning of GRB 230204B ’s radio data within this distribution provides valuable context for its radio emission profile relative to the larger GRB population. This comparison indicates that GRB 230204B ’s intrinsic radio luminosity is consistent with the broader sample, suggesting it shares similar radio emission characteristics with typical GRBs.
5 SUMMARY
In this study, we conducted a comprehensive analysis of GRB 230204B, a long GRB observed across multiple wavelengths. Our observational campaign utilized both space-based telescopes, including Fermi GBM, MAXI, Swift/BAT, Swift/XRT, and Swift/UVOT, as well as ground-based observatories like BOOTES, MASTER, DOT, and ATCA. The burst displayed a multi-pulse structure (each separated by periods of quiescence) with significant spectral variability, which we explored using time-resolved and time-integrated spectral analyses. Our findings indicate that the Band function, when combined with a Blackbody component, provided the best fit to the time-integrated spectra (indicating a potential mix of photospheric and synchrotron emission mechanisms), with variations in the preferred models across individual pulses. Furthermore, the time-resolved analysis reveals that the initial pulses of the prompt emission show harder spectra, while subsequent pulses become progressively softer, indicative of a hybrid jet composition.
Furthermore, the afterglow data collected across radio, optical, and X-ray wavelengths showed that GRB 230204B’s luminosity at various stages aligns with typical long-duration GRBs. Optical follow-up using multiple ground-based observatories revealed a rapid, early afterglow decay rate, which was more luminous than most GRBs at a similar epoch. Its high optical luminosity and bright afterglow suggest efficient energy transfer from the burst to the surrounding medium, or an environment conducive to high luminosity, possibly due to a structured jet.
Broadband afterglow modeling was conducted to investigate the GRB’s jet structure and surrounding environment. Using afterglowpy/Redback, we evaluated different jet profiles, concluding that a Top-Hat jet model best describes the data, with the inferred viewing and core angles suggesting a narrow, well-collimated jet structure. GRB 230204B exhibits an exceptionally high optical luminosity and an early steep optical decay, consistent with an observer viewing the burst nearly on-axis, as supported by afterglow modeling and a narrow inferred viewing angle. Our afterglow modeling supports a forward shock emission in a homogeneous interstellar medium (ISM), with no evidence of central engine activity. The optical luminosity and temporal decay index were found to align with existing GRB luminosity-decay correlations, further supporting a forward-shock-dominated afterglow. The radiative efficiency derived for GRB 230204B is lower ( 4.3%) compared to estimates reported for typical GRBs. The lower efficiency is primarily due to a lower magnetic field energy fraction, which leads to systematically higher estimates of . Consequently, this significantly mitigates the long-standing issue of low efficiency associated with internal shock models. We noted that GRB 230204B has a higher and values than expected from magnetar central engine (2 1052 erg). This supports the rapidly rotating black hole as a possible central engine for this energetic and bright burst.
In summary, GRB 230204B’s low gamma-ray efficiency and hybrid jet composition suggest a radiation mechanism where photospheric and synchrotron processes co-exist within the jet, influencing the spectral evolution of the prompt emission. This case highlights how the efficiency of GRB emissions depends not only on the jet composition and radiation processes but also on the structure of the jet and its interaction with the surrounding environment. Further studies of GRBs with similar spectral evolution and energy efficiency will help clarify the diversity of radiation mechanisms in GRBs and their implications for jet physics and energy dissipation.
Overall, our analysis of GRB 230204B underscores its importance as a reference for studying the emission mechanisms, jet structures, and afterglow behaviors of long-duration GRBs. The findings contribute to the growing understanding of GRB diversity and emphasize the role of broadband follow-up observations in exploring the complex evolution of these powerful cosmic events. Future studies of similar events, using well-coordinated observations across the spectrum, will be crucial in further refining our understanding of GRB physics and jet composition.
Appendix A Figures
| Tstart | Tstop | Flux | DICBand | Flux | DICCPL | DIC | |||||
| (sec) | (sec) | (keV) | (keV) | ||||||||
| -0.8 | 1.234 | 1.6 | 2080.12 | 1.31 | 2046.1 | 34.02 | |||||
| 1.234 | 4.973 | 2.54 | 2812.46 | 1.76 | 2812.03 | 0.43 | |||||
| 41.412 | 45.259 | 1.93 | 2844.64 | 1.78 | 2828.86 | 15.78 | |||||
| 45.259 | 60.297 | 3.52 | 4361.92 | 3.03 | 4369.86 | -7.94 | |||||
| 60.297 | 63.028 | 1.22 | 2501.94 | 1.09 | 2502.2 | -0.26 | |||||
| 63.028 | 68.152 | 0.63 | 3138.76 | 0.42 | 3131.87 | 6.89 | |||||
| 109.451 | 122.927 | 0.71 | 4189.15 | 0.64 | 4172.96 | 16.19 | |||||
| 122.927 | 140.987 | 1.9 | 4522.29 | 1.79 | 4523.35 | -1.06 | |||||
| 140.987 | 155.493 | 1.56 | 4261.59 | 1.47 | 4264.6 | -3.01 | |||||
| 155.493 | 156.992 | 2.22 | 1898.35 | 2.01 | 1900.76 | -2.41 | |||||
| 172.583 | 190.95 | 0.99 | 4436.83 | 0.94 | 4443.64 | -6.81 | |||||
| 192.363 | 198.11 | 0.63 | 3236.22 | 0.57 | 3242.3 | -6.08 | |||||
| 198.11 | 200.125 | 0.82 | 2116.74 | 0.77 | 2114.43 | 2.31 | |||||
| 200.125 | 230.0 | 0.39 | 4862.37 | 0.37 | 4862.75 | -0.38 |
| Time-T0 (sec) | Energy (keV) | FD (uJy) | FD Err | Telescope | Ref. |
| 91528.12 | 10.0 | 0.0116 | 0.0030 | XRT | This work |
| 848730.52 | 10.0 | 0.0038 | 0.0019 | XRT | This work |
| Time-T0 (sec) | Filter | MAG (AB) | MAG Err | Telescope | Ref. |
| 1378.42698 | unfiltered | 12.7 | 0.12 | MASTER | This work |
| 1777.40663 | unfiltered | 12.87 | 0.11 | MASTER | This work |
| 2930.5762 | unfiltered | 13.77 | 0.17 | MASTER | This work |
| 1389.55696 | unfiltered | 12.74 | 0.1 | MASTER | This work |
| 1611.76601 | unfiltered | 13.03 | 0.1 | MASTER | This work |
| 1830.95002 | unfiltered | 13.18 | 0.1 | MASTER | This work |
| 2048.95398 | unfiltered | 13.4 | 0.11 | MASTER | This work |
| 2267.91196 | unfiltered | 13.41 | 0.08 | MASTER | This work |
| 2486.30301 | unfiltered | 13.54 | 0.07 | MASTER | This work |
| 2694.05396 | unfiltered | 13.6 | 0.07 | MASTER | This work |
| 2914.54193 | unfiltered | 13.74 | 0.08 | MASTER | This work |
| 3654.57493 | unfiltered | 14.2 | 0.07 | MASTER | This work |
| 3877.80896 | unfiltered | 14.23 | 0.04 | MASTER | This work |
| 4097.35897 | unfiltered | 14.36 | 0.04 | MASTER | This work |
| 4316.58904 | unfiltered | 14.41 | 0.04 | MASTER | This work |
| 2296.5 | clear | 13.51 | 0.06 | BOOTES | This work |
| 2300.8 | clear | 13.4 | 0.29 | BOOTES | This work |
| 2319.4 | clear | 13.5 | 0.13 | BOOTES | This work |
| 2345.0 | clear | 13.37 | 0.13 | BOOTES | This work |
| 2363.8 | clear | 13.33 | 0.17 | BOOTES | This work |
| 2369.8 | clear | 13.65 | 0.16 | BOOTES | This work |
| 2375.9 | clear | 13.29 | 0.16 | BOOTES | This work |
| 2382.0 | clear | 13.66 | 0.14 | BOOTES | This work |
| 2399.4 | clear | 13.65 | 0.14 | BOOTES | This work |
| 2421.5 | clear | 13.7 | 0.15 | BOOTES | This work |
| 2427.6 | clear | 13.52 | 0.13 | BOOTES | This work |
| 2433.9 | clear | 13.5 | 0.11 | BOOTES | This work |
| 2439.9 | clear | 13.55 | 0.28 | BOOTES | This work |
| 2446.0 | Clear | 13.53 | 0.21 | BOOTES | This work |
| 2459.1 | Clear | 13.4 | 0.1 | BOOTES | This work |
| 2470.6 | Clear | 13.55 | 0.19 | BOOTES | This work |
| 2491.9 | Clear | 13.62 | 0.2 | BOOTES | This work |
| 2517.2 | Clear | 13.54 | 0.15 | BOOTES | This work |
| 2523.4 | Clear | 13.55 | 0.18 | BOOTES | This work |
| 2545.4 | Clear | 13.65 | 0.14 | BOOTES | This work |
| 3269.6 | Clear | 14.02 | 0.06 | BOOTES | This work |
| 3308.6 | Clear | 13.98 | 0.14 | BOOTES | This work |
| 3335.6 | Clear | 14.1 | 0.08 | BOOTES | This work |
| 3360.6 | Clear | 14.14 | 0.14 | BOOTES | This work |
| 3396.6 | Clear | 14.18 | 0.08 | BOOTES | This work |
| 3421.6 | Clear | 14.13 | 0.06 | BOOTES | This work |
| 3446.6 | Clear | 14.1 | 0.09 | BOOTES | This work |
| 3472.6 | Clear | 14.18 | 0.06 | BOOTES | This work |
| 3497.6 | Clear | 14.2 | 0.08 | BOOTES | This work |
| 3523.6 | Clear | 14.15 | 0.09 | BOOTES | This work |
| 3548.6 | Clear | 14.19 | 0.04 | BOOTES | This work |
| 3574.6 | Clear | 14.21 | 0.09 | BOOTES | This work |
| 3599.6 | Clear | 14.16 | 0.08 | BOOTES | This work |
| 3626.6 | Clear | 14.25 | 0.08 | BOOTES | This work |
| 3651.6 | Clear | 14.23 | 0.08 | BOOTES | This work |
| 3677.6 | Clear | 14.27 | 0.13 | BOOTES | This work |
| 3703.6 | Clear | 14.25 | 0.08 | BOOTES | This work |
| 3729.6 | Clear | 14.3 | 0.07 | BOOTES | This work |
| 3755.6 | Clear | 14.38 | 0.09 | BOOTES | This work |
| 3781.6 | Clear | 14.35 | 0.13 | BOOTES | This work |
| 3806.6 | Clear | 14.35 | 0.08 | BOOTES | This work |
| 3831.6 | Clear | 14.33 | 0.05 | BOOTES | This work |
| 3858.6 | Clear | 14.38 | 0.07 | BOOTES | This work |
| 3884.6 | Clear | 14.4 | 0.1 | BOOTES | This work |
| 3909.6 | Clear | 14.44 | 0.06 | BOOTES | This work |
| 3935.6 | Clear | 14.44 | 0.09 | BOOTES | This work |
| 3960.6 | Clear | 14.46 | 0.05 | BOOTES | This work |
| 3988.6 | Clear | 14.48 | 0.07 | BOOTES | This work |
| 4028.6 | Clear | 14.45 | 0.08 | BOOTES | This work |
| 101119.0 | 21.09 | 0.09 | DOT | This work | |
| 278325.4 | 21.27 | – | DOT | This work | |
| 103451.8 | 20.88 | 0.10 | DOT | This work | |
| 200738.2 | 21.67 | 0.12 | DOT | This work | |
| 105611.8 | 22.27 | – | DOT | This work | |
| 95569.7 | 18.94 | UVOT | This work | ||
| 95159.7 | 19.50 | UVOT | This work | ||
| 95075.4 | 20.57 | UVOT | This work | ||
| 93868.0 | 19.53 | UVOT | This work | ||
| 89647.6 | 21.7 | UVOT | This work | ||
| 89126.7 | 19.78 | UVOT | This work | ||
| 841905.2 | 20.81 | UVOT | This work | ||
| 5889.9 | 15.25 | 0.04 | GIT | GCN | |
| 8801.5 | 16.20 | 0.05 | GIT | GCN | |
| 12905.8 | 16.82 | 0.04 | ATLAS | GCN | |
| 13064.8 | 16.95 | 0.05 | ATLAS | GCN | |
| 14204.8 | 17.10 | 0.05 | ATLAS | GCN | |
| 14477.8 | 17.09 | 0.05 | ATLAS | GCN | |
| 110948.8 | 21.30 | 0.18 | VLT | GCN |
| Time-T0 (sec) | Freq (GHz) | FD (mJy) | FD Err | Telescope | Ref. |
| 314133 | 5.5 | 0.45 | ATCA | This work | |
| 314133 | 9.0 | 0.20 | ATCA | This work | |
| 314133 | 16.7 | 0.21 | 0.02 | ATCA | This work |
| 314133 | 21.2 | 0.24 | 0.05 | ATCA | This work |
| 314133 | 33.0 | 0.12 | ATCA | This work | |
| 314133 | 35.0 | 0.13 | ATCA | This work | |
| 12658049 | 0.842 | 1.31 | ASKAP | This work | |
| 28948293 | 0.943 | 1.45 | ASKAP | This work |
| T-T0 (sec) | Energy range exposure | Upper limit (mJy) |
| 22526 | 2-20 keV | 0.052 |
| 28102 | 2-20 keV | 0.042 |
| 33679 | 2-20 keV | 0.047 |
| 39257 | 2-20 keV | 0.047 |
| 44834 | 2-20 keV | 0.058 |
| 50409 | 2-20 keV | 0.062 |
| 55986 | 2-20 keV | 0.058 |
| 61563 | 2-20 keV | 0.057 |
| 67140 | 2-20 keV | 0.053 |
| 72717 | 2-20 keV | 0.055 |
| 78292 | 2-20 keV | 0.055 |

### v2

@@
-Extremely luminous optical afterglow of a distant and energetic gamma-ray burst GRB 230204B
+Extremely luminous optical afterglow of an energetic gamma-ray burst GRB 230204B
 Abstract
@@
 Time-resolved spectral analysis reveals a transition in the prompt emission from hard (sub-photospheric dominated) spectra during early pulses to softer (synchrotron radiation dominated) spectra in later pulses, indicative of a hybrid jet composition.
-We report the discovery and characterization of the optical afterglow using the MASTER and BOOTES robotic telescope networks, alongside long-term radio observations extending to 335 days post-burst with the ATCA.
-At 1.3 ks post-burst, the optical luminosity was exceptionally high, surpassing even other bright GRBs, such as GRB 221009A (the “BOAT”).
+We report the discovery and characterization of the optical afterglow using the MASTER and BOOTES robotic telescope networks, which enabled rapid follow-up observations starting at 1.3 ks post-burst.
+The optical luminosity at this time was exceptionally high, surpassing that of many other optically bright GRBs, such as GRB 990123, GRB 080319B, etc.
+This places the burst among the most luminous optical GRBs observed to date.
+Long-term radio observations extending to 335 days post-burst were conducted with the ATCA.
 Multi-wavelength modeling, incorporating data from MASTER, BOOTES, DOT, Swift/XRT, and radio observations, was conducted using an external ISM forward-shock top-hat jet model with afterglowpy.
@@
 1 Introduction
-Gamma-ray bursts (GRBs) are among the most energetic events in the Universe, emitting vast amounts of energy within a few seconds to minutes, often followed by an afterglow that can be observed across the electromagnetic spectrum (Piran, 1999; Kumar & Zhang, 2015).
-The discovery of afterglows in the late 1990s marked a significant milestone in GRB research, enabling precise localization and redshift determination, which, in turn, provided insights into the distances and energies involved (Costa et al., 1997; van Paradijs et al., 1997; Frail et al., 1997).
-The study of GRB afterglows, particularly in the optical regime, has been pivotal in understanding the progenitor systems, the environment surrounding the burst, and the mechanisms driving the emission (Kann et al., 2010; Li et al., 2012; Gupta, 2023).
-The temporal evolution of these afterglows, often characterized by a power-law decay, can reveal crucial details about the jet dynamics, energy injection, and interaction with the surrounding medium (Meszaros & Rees, 1997a; Sari et al., 1998).
-In recent times, robotic telescope networks, such as Mobile Astronomical System of Telescope-Robots (MASTER), Burst Observer and Optical Transient Exploring System (BOOTES), and others, have become indispensable in the rapid follow-up of GRBs, often identifying optical afterglows within minutes to hours (especially for Fermi Gamma-ray Burst Monitor (GBM) detected GRBs with large error circles) of the initial burst (Lipunov et al., 2010a; Castro-Tirado et al., 2012).
-Fermi GBM’s localization accuracy, while adequate for high-energy studies, often leaves large localization uncertainties spanning a few to tens of degrees radius, making rapid and wide-field follow-up observations essential.
+Gamma-ray bursts (GRBs) are among the most energetic events in the Universe, emitting vast amounts of energy within a few seconds to minutes, often followed by an afterglow that can be observed across the electromagnetic spectrum (1995ARA&A..33..415F; 1997ApJ...476..232M; 1998ApJ...497L..17S; 1999PhR...314..575P; 2000ApJ...543...66P; 2001ApJ...562L..55F; 2004RvMP...76.1143P; 2006RPPh...69.2259M; 2006ApJ...642..354Z; 2006ARA&A..44..507W; 2014ARA&A..52...43B; 2015PhR...561....1K; 2017ApJ...848L..12A).
+The discovery of afterglows in the late 1990s marked a significant milestone in GRB research.
+The first detection of an X-ray afterglow by 1997Natur.387..783C, optical afterglow by 1997Natur.386..686V, and radio afterglow by 1997Natur.389..261F enabled precise localization and redshift determination.
+This, in turn, provided insights into the distances and energies involved in GRBs (1999Natur.398..389K).
+The study of GRB afterglows, particularly in the optical regime, has been pivotal in understanding the progenitor systems (2001Sci...291...79M), the environment surrounding the burst (2010ApJ...720.1513K; 2012ApJ...758...27L), and the mechanisms driving the emission (2023arXiv231216265G).
+The temporal evolution of these afterglows, often characterized by a power-law decay, can reveal crucial details about the jet dynamics, energy injection, and interaction with the surrounding medium (1997ApJ...476..232M; 1998ApJ...497L..17S).
+In recent times, robotic telescope networks, such as Mobile Astronomical System of Telescope-Robots (MASTER, Li2022), Burst Observer and Optical Transient Exploring System (BOOTES, 2023NatAs...7.1136C), Las Cumbres Observatory (LCO) Global Telescope (2013PASP..125.1031B), Rapid Eye Mount (REM, 2004SPIE.5492.1613C), Liverpool Telescope (LT, 2016RMxAC..48...48C), Gravitational-wave Optical Transient Observer (GOTO, 2022MNRAS.511.2405S), Multicolor Imaging Telescopes for Survey and Monstrous Explosions (MITSuME, 2008AIPC.1000..543S), the Robotic Optical Transient Search Experiment (ROTSE III, 2003PASP..115..132A), Nazarbayev University Transient Telescope at Assy-Turgen Astrophysical Observatory (NUTTelA-TAO, 2022SPIE12184E..8AG), and others, have become indispensable in the rapid follow-up of GRBs, often identifying optical afterglows within minutes to hours (especially for Fermi Gamma-ray Burst Monitor (GBM) detected GRBs with large error circles) of the initial burst (2010AdAst2010E..30L; 2012ASInC...7..313C).
+Fermi GBM’s localization accuracy, while adequate for high-energy studies, often leaves large localization uncertainties spanning a few to tens of degrees radius (2009ApJ...702..791M; 2015ApJS..216...32C; 2020ApJ...895...40G), making rapid and wide-field follow-up observations essential.
 These robotic telescope networks, equipped with wide-field cameras and fully automated systems, are specifically designed to address this challenge.
 Their ability to quickly respond to GRB alerts and scan large areas of the sky allows them to pinpoint transient optical emissions within the Fermi error regions.
-Additionally, their prompt response times enable them to observe early afterglow phases, providing data at these early times critical for informing the onset and evolution of the emission (Gupta et al., 2021b; Ror et al., 2023b; Sánchez-Ramírez et al., 2024).
-While these observations are useful for constraining GRB redshifts indirectly by identifying afterglow candidates for follow-up spectroscopic studies, they typically rely on single-band detections and are limited in sensitivity, restricting their ability to detect only the bright afterglows.
-Despite these limitations, the robotic nature of these observatories ensures continuous and autonomous temporal coverage, improving the chances of detecting GRBs in real-time and providing crucial data to guide further multi-wavelength follow-up efforts (Castro-Tirado, 2023).
-GRB 230204B, detected by the Fermi GBM and MAXI, stands out as one of the most energetic GRBs observed, with an isotropic equivalent energy (Eγ,iso) exceeding 1054 erg.
-The optical afterglow of this burst was discovered by the MASTER and BOOTES robotic telescope networks approximately 1.3 ks post-burst, marking the beginning of the broadband follow-up campaign.
-Additionally, at the time of MASTER/BOOTES detection ( 1.3 ks post-burst), the optical luminosity was found to be exceptionally high when compared to other bright bursts, suggesting that GRB 230204B harbored an extremely luminous afterglow phase.
-In this paper, we present a detailed analysis of the optical and multi-wavelength follow-up observations (see Section 2) of GRB 230204B, including data from MASTER, BOOTES, Devasthal Optical Telescope (DOT), Swift Ultraviolet and Optical telescope (UVOT) and radio telescope.
+Additionally, their prompt response times enable them to observe early afterglow phases, providing data at these early times critical for informing the onset and evolution of the emission (2021MNRAS.505.4086G; 2023ApJ...942...34R; 2024A&A...692A...3S).
+These observations are useful for constraining GRB redshifts indirectly by identifying afterglow candidates for follow-up spectroscopic studies.
+Furthermore, the robotic nature of these observatories ensures continuous and autonomous temporal coverage, improving the chances of detecting GRBs in real-time and providing crucial data to guide further multi-wavelength follow-up efforts (2023NatAs...7.1136C).
+GRB 230204B, detected by the Fermi GBM and MAXI, stands out as one of the most energetic GRBs observed, with an isotropic equivalent energy (Eγ,iso) exceeding 1054 erg (see Section 3.2).
+The rapid and precise localization of GRB 230204B by the MAXI detector triggered an immediate response from the MASTER and BOOTES robotic telescope networks, which provided crucial early-time optical data that significantly shaped our physical interpretation of this burst.
+These networks are specifically designed for the rapid follow-up of transient astronomical events, such as GRBs, and their automated observing capabilities are crucial for observing the early phases of the optical afterglow (2023NatAs...7.1136C).
+The MASTER network began observing GRB 230204B at 1.3 ks post-burst, providing the earliest optical detection of the afterglow of GRB 230204B and constrained the exceptionally high optical luminosity at the time of MASTER/BOOTES detection ( 1.3 ks post-burst), allowing us to place GRB 230204B among the most luminous optical afterglows known.
+The BOOTES network complemented these observations with additional early-time data, which were essential for constraining the decay phase of the optical light curve (see Section 3.3).
+In this paper, we present a detailed analysis of the optical and multi-wavelength follow-up observations (see Section 2) of GRB 230204B, including data from MASTER (2023GCN.33441....1L), BOOTES (2023NatAs...7.1136C), Devasthal Optical Telescope (DOT, 2023GCN.33284....1R; 2024BSRSL..93..683G), Swift Ultraviolet and Optical telescope (UVOT, 2023GCN.33292....1S; 2023Univ....9..113O), Growth India Telescope (GIT, 2023GCN.33269....1S), Asteroid Terrestrial-impact Last Alert System (ATLAS, 2023GCN.33278....1S), Very Large Telescope (VLT, 2023GCN.33281....1S), Australia Telescope Compact Array (ATCA, 2023GCN.33321....1G), and Australian Square Kilometre Array Pathfinder (ASKAP, 2020PASA...37...48M) telescope.
 We explore the temporal and spectral characteristics of the prompt emission and afterglow and compare its optical luminosity with other known GRBs, the results are presented in Section 3.
@@
 2 Observations and data analysis
-This section details the observation campaign for GRB 230204B, observed by multiple space-based and ground-based telescopes across a broad wavelength range, from gamma rays to radio frequencies.
-Comprehensive details of the observing campaign are provided below and the sequence of prompt and afterglow observations for GRB 230204B is shown in Figure 1.
+This section details the observation campaign for GRB 230204B, observed by multiple space-based and ground-based telescopes across a broad wavelength range.
+Comprehensive details of the observing campaign are provided below, and the sequence of prompt and afterglow observations for GRB 230204B is shown in Figure 1.
 2.1 Prompt Emission
-GRB 230204B was first detected on 4th February 2023 by the Fermi satellite at 21:44:27.20 UT (hereafter T0), with duration of 216 seconds in 50-300 keV (Poolakkil et al., 2023).
-MAXI also detected it at 21:47:51 UT and the burst position was constrained to approximately RA = 13h 10m 19s, Dec = -21° 45′ 07′′ (J2000) with a statistical uncertainty region of 0.1°(Serino et al., 2023).
-The burst did not trigger the Neil Gehrels Swift Observatory; henceforth Swift (Gehrels et al., 2004).
-However, following the MAXI trigger notification, the Gamma-ray Urgent Archiver for Novel Opportunities (GUANO, Tohuvavohu et al.
-2020) pipeline requested data from the onboard buffer of the Swift Burst Alert Telescope (BAT) and initiated a search within the 200-second event mode data centered around T0.
-The GRB was detected in this Swift/BAT GUANO data, determining its to be approximately 65 seconds (Kennea et al., 2023).
-The burst was also detected by the Indian multiwavelength satellite AstroSat, which measured a of seconds (Waratkar et al., 2023).
-The prompt emission of the burst was also seen by AGILE (Casentini et al., 2023) and GRBAlpha (Dafcikova et al., 2023) missions, with every mission providing a slightly different value of duration as expected due to the different sensitive and spectral coverage.
+GRB 230204B was first detected on 4th February 2023 by the Fermi GBM satellite at 21:44:27.20 UT (hereafter T0), with duration of 216 seconds in 50-300 keV (2023GCN.33288....1P).
+MAXI also detected it at 21:47:51 UT and the burst position was constrained to approximately RA = 13h 10m 19s, Dec = -21° 45′ 07′′ (J2000) with a statistical uncertainty region of 0.1°(2023GCN.33265....1S).
+The burst did not trigger the Neil Gehrels Swift Observatory; henceforth Swift (2004ApJ...611.1005G).
+However, following the MAXI trigger notification, the Gamma-ray Urgent Archiver for Novel Opportunities (GUANO, 2020ApJ...900...35T) pipeline requested data from the onboard buffer of the Swift Burst Alert Telescope (BAT) and initiated a search within the 200-second event mode data centered around T0.
+The GRB was detected in this Swift/BAT GUANO data, determining its to be approximately 65 seconds (2023GCN.33267....1K).
+The burst was also detected by the Indian multiwavelength satellite AstroSat, which measured a of seconds (2023GCN.33268....1W).
+The prompt emission of the burst was also seen by AGILE duration 145 sec, 2023GCN.33272....1C) and GRBAlpha ( of 207 sec, 2023GCN.33273....1D) missions, with every mission providing a slightly different value of duration as expected due to the different sensitive and spectral coverage.
 2.2 Prompt X-ray Emission: MAXI
-MAXI (mounted on the ISS) scans a position of the sky every 92 minutes (Matsuoka et al., 2009).
+MAXI (mounted on the ISS) scans a position of the sky every 92 minutes (maxi).
 The position of the GRB moved into the MAXI field of view at T0 +192.8 sec, and the scan transit ended at T0 +248.8 sec.
-During the scan transit, the MAXI/GSC nova alert system (Negoro et al., 2016) triggered the burst at 21:47:51 UT (T0 + 203.8 sec), with MAXI’s detection revealing the soft nature of the burst.
-Assuming the source flux was constant over the transit, the X-ray flux in the 4.0–10.0 keV band was calculated to be 628 49 mCrab (1 sigma error, Serino et al.
-2023).
-After the detection of prompt emission, MAXI/GSC could not observe the position of the burst for about 6 hours due to the observation constraint of the radiation zone (Sugizaki et al., 2011) in the International Space Station (ISS) orbit.
-The observation resumed at T0 + 22.5 ks, and there were 11 consecutive scan transits before the start of the observations by Swift/XRT.
-The afterglow was not detected in any transit.
-We calculated 3 sigma upper limits in the 2–20 keV band (see Table 8 in appendix) for these transits following the method by Sugita et al.
-(2018).
+During the scan transit, the MAXI/GSC nova alert system (negoro) triggered the burst at 21:47:51 UT (T0 + 203.8 sec), with MAXI’s detection revealing the soft nature of the burst.
+The detection occurred during the final stages of the GRB emission, as evidenced by its temporal proximity to the main burst and the significant variability observed in the MAXI light curve during the scan transit (see Figure 4).
+The observed temporal variability is characteristic signature of prompt emission (2004IJMPA..19.2385Z; 2015PhR...561....1K; 2023A&A...671A.112C) rather than early afterglow, which typically exhibits smoother temporal evolution (1998ApJ...497L..17S; 2006ApJ...642..389N; 2006ApJ...647.1213O).
+These characteristics support the interpretation that the detected X-rays are part of the prompt emission of GRB 230204B rather than the afterglow.
+Assuming the source flux was constant over the transit, the X-ray flux in the 4.0–10.0 keV band was calculated to be 628 49 mCrab (1 sigma error, 2023GCN.33265....1S).
+After the detection of the late phase of prompt emission of GRB 230204B by MAXI/GSC, it could not observe the position of the burst for about 6 hours due to the observation constraint of the radiation zone (sugizaki) in the International Space Station (ISS) orbit.
+The MAXI observations resumed at T0 + 22.5 ks, and there were 11 consecutive scan transits to search for an X-ray afterglow of GRB 230204B.
+The afterglow was not detected in any transit of MAXI observations.
+We calculated 3 sigma upper limits in the 2–20 keV band (see Table 9 in appendix) for these transits following the method by sugita.
+This non-detection by MAXI/GSC during the late phase further supports the prompt nature of the initial MAXI detection.
+A deeper search for the X-ray counterpart of GRB 230204B was later ( T0 + 80 ks) carried out by Swift/XRT (see Section 2.3).
 2.3 Afterglow: Soft X-ray Emission
-The fading X-ray counterpart of GRB 230204B was detected by Swift/XRT approximately 80 ks after the burst trigger (D’Elia et al., 2023).
-Given the faint nature of the X-ray afterglow at such a late epoch, all observations were performed in photon counting (PC) mode to optimize sensitivity for faint sources.
+The fading X-ray counterpart of GRB 230204B was detected by Swift/XRT approximately 80 ks after the burst trigger (2023GCN.33285....1D).
+Given the faint nature of the X-ray afterglow at such a late epoch, all observations were performed in photon counting (PC) mode to optimize sensitivity for faint sources (2007A&A...469..379E; 2009MNRAS.397.1177E).
 Follow-up observations were conducted during two epochs (XRT exposure times of 5.1 ks, and 4.5 ks, respectively), on 5 February 2023 and 14 February 2023, resulting in only two usable data points (see Table A2 of the appendix).
@@
 We discovered the optical afterglow of GRB 230204B using two robotic telescope networks (using MASTER in the unfiltered band and using BOOTES in the clear band, respectively) independently (see Figure 2).
-The optical counterpart of the burst was also identified by the 0.7m GROWTH-India Telescope (GIT) approximately 1.57 hours (later than MASTER and BOOTES) after the burst trigger (Swain et al., 2023).
-This optical afterglow was subsequently observed by several other ground-based telescopes (Smartt et al., 2023; Ror et al., 2023a).
-The redshift of the burst, = 2.142, was determined through spectroscopic observations using the X-shooter spectrograph on the Very Large Telescope (VLT, Saccardi et al.
-2023).
+The optical counterpart of the burst was also identified by the 0.7m GROWTH-India Telescope (GIT) approximately 1.57 hours (later than MASTER and BOOTES) after the burst trigger (2023GCN.33269....1S).
+This optical afterglow was subsequently observed by several other ground-based telescopes (2023GCN.33278....1S; 2023GCN.33284....1R).
+Using the X-shooter spectrograph on the Very Large Telescope (VLT), 2023GCN.33281....1S estimated a redshift of = 2.142 for the burst.
 The given magnitude (for the positive detection) and upper limits within 3 are listed in Table LABEL:tab:optical.
 2.4.1 MASTER robotic telescope
-MASTER global network consists of 9 identical fully robotic wide-field twin telescopes with 4-8 square degrees field of view, distributed around the Earth for all-sky monitoring up to 19-20 mag with the goal of detecting optical transients, including high energy astrophysics events such as GRBs, Gravitational waves, ultra-high energy neutrino and Fast Radio Bursts (Lipunov et al., 2010b, 2019, 2022, 2023).
-We observed such events as soon as possible upon receiving the alerts (Ershova et al., 2020) using the GCN system (Barthelmy et al., 1998; Barthelmy, 2003, 2016) every night in both, the fully robotic open (8 square degrees FOV) and closed (4 square degrees FOV) modes, being ready for observations with fully clear hemisphere (clamshell dome constructions).
-Open mode is used in the case of alerts provided by Fermi, LIGO/Virgo, IceCube, ANTARES/KM3Net (with large error boxes) when at least 2 images per night are required if large error boxes are at least 5 degrees above the horizon, in order to cover as much as possible with our given field of view.
-Closed mode (using both tubes in parallel Kornilov et al.
-(2012)) is preferred for events with error boxes smaller than 1 degree such as the ones provided by Swift, MAXI, Fermi-LAT, GECAM, Einstein Probe, INTEGRAL, when polarization observations can be undertaken (Troja et al., 2017; Jordana-Mitjans et al., 2020; Gupta et al., 2022a).
-The first exposure duration also depends on the time between the trigger and the start of observation and can be increased from 10 sec at the first minute to 180 sec at later stages in order to detect the fainter objects.
-GRB 230204B was observed by MASTER-Kislovodsk and MASTER-SAAO.
+The MASTER global network comprises nine identical fully robotic wide-field twin telescopes, distributed around the Earth for all-sky monitoring up to 19-20 mag.
+Its primary goal is to detect optical transients, including high-energy astrophysical events such as GRBs (Li2010; Li2019; Li2022; Li2023).
+The network operates in both open (8 square degrees FOV) and closed (4 square degrees FOV) modes, enabling rapid follow-up observations of alerts from facilities such as Fermi, LIGO/Virgo, IceCube, ANTARES/KM3Net, Swift, MAXI, Fermi-LAT, GECAM, Einstein Probe, and INTEGRAL (Kornilov; Troja; Mundell; Gupta1).
+GRB 230204B was observed by MASTER-Kislovodsk and MASTER-SAAO (South African Astronomical Observatory).
 First images were received for the MAXI error box, then telescopes automatically pointed to Fermi field.
 The MASTER-Kislovodsk robotic telescope started observing the MAXI trigger 979708795 observation at 2023-02-04 22:05:55 UT (1064 sec after trigger time which was 74 sec after the notification time, see Figure 2) at 11∘ altitude, in 44∘ distance to the Moon (0.3 phase, that creates significant interference for wide-field instruments observations) through clouds at horizon with =14.8 mag at first image in polarization filter.
+The transient was not detected in this initial image.
 The MASTER-SAAO robotic telescope located in the South African Astronomical Observatory, started observing error box automatically 85 seconds after notice time (i.e.
@@
 The photometry dataset excludes frames where the comparison star data were deemed unreliable.
-The magnitude error for comparison stars was computed as the deviation between the weighted average magnitude (where weights are the inverse square of the signal-to-noise ratio) and the magnitude measured on individual frames.
+The magnitude errors for comparison stars were computed as the deviation between the weighted average magnitude (where weights are the inverse square of the signal-to-noise ratio) and the magnitude measured on individual frames.
 Two approaches were employed to determine the stellar magnitude.
 In the first approach, a reference star from Gaia EDR3 was identified, its flux on the observational frame measured, and the magnitude calculated using Pogson’s scale.
-In the second approach, a zero-point correction was applied to the reference star’s magnitude, utilizing the comparison stars as secondary references.
+This method served as a validation step to assess consistency and identify potential systematic errors.
+In the second approach, which was used for the final magnitude determination, a zero-point correction was applied to the reference star’s magnitude, utilizing the comparison stars as secondary references.
 The second method generally produced a smaller magnitude variance compared to the first, owing to the improved calibration achieved through the zero-point correction.
-This dual-approach methodology ensures robust photometric accuracy while mitigating systematic errors.
+By employing this dual-approach methodology, we ensured robust photometric accuracy while mitigating systematic errors.
+The final results presented in this study are based solely on the second method.
 2.4.2 BOOTES robotic telescope
@@
 The BOOTES-2/TELMA telescope and the BOOTES-5/JGT telescope also executed follow-up observations at 2023-02-05 04:16:55 UT and 2023-02-05 07:43:45 UT, respectively.
-A series of images were obtained using the clear filter with exposures of 1 sec, 5 sec, and 10 sec at BOOTES-4/MET telescope during the first epoch and the optical afterglow was clearly detected during the decaying phase.
-However, during the following multi-filters (Sloan- and clear filters, 60 sec exposure) monitoring, the optical afterglow was not detected on the stacked images for each epoch (as they were not as deep as the clear filter exposures).
-All the images were processed with bias, dark subtraction, and flat-field correction using custom IRAF routines (Gupta et al., 2021b).
-Photometry was carried out using the standard IRAF package and the images were calibrated with nearby comparison stars from the USNO-B1.0 and SDSS catalogs.
+A series of images were obtained using the clear filter with exposures of 1 sec, 5 sec, and 10 sec at BOOTES-4/MET telescope during the first epoch, and the optical afterglow was clearly detected during the decaying phase.
+However, during the subsequent multi-filter monitoring (Sloan- and clear filters, 60 sec exposure), the optical afterglow was not detected on the stacked images for each epoch.
+This non-detection is likely due to the fact that the late-time stacked images were not deep enough to detect the rapidly fading afterglow.
+All the images were processed with bias, dark subtraction, and flat-field correction using custom IRAF routines (2021MNRAS.505.4086G).
+Photometry was carried out using the standard IRAF package, and the images were calibrated with nearby comparison stars from the USNO-B1.0 (for clear filter) and SDSS (for filters) catalogs.
 2.4.3 DOT
-We initiated observations using the DOT (Pandey, 2016; Gupta et al., 2024a) on 2023-02-06 at 01:49:47.939 UT, approximately 1.16 days after the trigger.
+We initiated observations using the DOT (2016RMxAC..48...83P; 2024BSRSL..93..683G) on 2023-02-06 at 01:49:47.939 UT, approximately 1.16 days after the trigger.
 Multiple frames with an exposure time of 180 sec each were observed in the , , and bands, and continuous observations were conducted over four consecutive nights using the same filters.
@@
 We stacked the images for each night in the same filter to increase the signal-to-noise ratio.
-PSF photometry was performed on the final stacked images (Gupta et al., 2022b, c).
-Due to the high moon phase during the burst observation, we only detected GRB 230204B in the and -band images from the first night.
+PSF photometry was performed on the final stacked images (2022JApA...43...82G; 2022JApA...43...11G).
+Due to the bright moon phase during the burst observation, we only detected GRB 230204B in the and -band images from the first night.
 For the remaining filters, we could only establish upper limits.
@@
 2.4.4 Swift/UVOT
-The Swift/UVOT began settled observations of the field of GRB 230204B 80.6 ks after the MAXI trigger (Serino et al., 2023).
+The Swift/UVOT began settled observations of the field of GRB 230204B 80.6 ks after the MAXI trigger (2023GCN.33265....1S).
 No optical afterglow was detected by UVOT.
@@
 Background counts were extracted using an aperture of 20 arcsec in radius located in a source-free region near the GRB.
-We use the Swift tool uvotsource to obtain the background subtracted count rates using source and background apertures of 5 and 20 arcseconds in radius, respectively.
-They were converted to magnitude using the UVOT photometric zero points (Breeveld et al., 2011; Poole et al., 2008).
+We use the Swift tool uvotsource to obtain the background-subtracted count rates using source and background apertures of 5 and 20 arcseconds in radius, respectively.
+They were converted to magnitude using the UVOT photometric zero points (bre11; poole).
 2.4.5 Radio Emission
-We observed GRB 230204B on 8 February 2023 (3.6 days post-burst), with the Australia Telescope Compact Array (ATCA, Project code CX527) at a wide range of frequencies centered on 5.5, 9.0, 16.7, 21.2, 33 and 35 GHz, each with a 2048 MHz-wide band.
-We reduced the visibility data using standard routines in MIRIAD (Sault et al., 1995).
+We observed GRB 230204B on 8 February 2023 (3.6 days post-burst), with the Australia Telescope Compact Array (ATCA, Project code CX527) at a wide range of frequencies centered on 5.5, 9.0, 16.7, 21.2, 33, and 35 GHz, each with a 2048 MHz-wide band.
+We reduced the visibility data using standard routines in MIRIAD (1995ASPC...77..433S).
 We used a combination of manual and automatic RFI flagging before calibration, conducted with MIRIAD tasks uvflag and pgflag, respectively.
-We used PKS 193463 to determine the bandpass response for frequency bands 5.5 and 9 GHz, and PKS 0727115 as the bandpass calibrator for 16.7, 21.2, 33 and 35 GHz bands.
+We used PKS 193463 to determine the bandpass response for frequency bands 5.5 and 9 GHz, and PKS 0727115 as the bandpass calibrator for 16.7, 21.2, 33, and 35 GHz bands.
 We used PKS 193463 to calibrate the flux density scale for all frequency bands.
 We used PKS 1256220 to calibrate the time-variable complex gains for all epochs and frequency bands.
-After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (Clark, 1980) on the target source field using standard MIRIAD tasks INVERT, CLEAN and RESTOR to obtain the final images.
-We detect the radio counterpart at 16.7 and 21.2 GHz at a position consistent with the BOOTES/MASTER/GIT optical counterpart position (Swain et al., 2023).
-For all other frequencies, we report 5-sigma upper limits (see Table 7 of appendix).
-The GRB position was also observed with the Australian SKA Pathfinder by the Rapid ASKAP Continuum Survey (RACS; McConnell et al., 2020) on 2023 July 1 and 2024 Jan 5.
-We find no radio detections in preliminary data and report 5-sigma upper limits of 1.5 mJy.
-There were also no archival radio detection pre-burst within 1 arcminute in surveys: the National Radio Astronomy Observatory VLA Sky Survey (Condon et al., 1998), the Sydney University Molonglo Sky Survey (Mauch et al., 2003), the Very Large Array Sky Survey (Lacy et al., 2020) or the RACS survey.
+After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (1980A&A....89..377C) on the target source field using standard MIRIAD tasks INVERT, CLEAN, and RESTOR to obtain the final images.
+We detect the radio counterpart at 16.7 and 21.2 GHz at a position consistent with the BOOTES/MASTER/GIT optical counterpart position (2023GCN.33269....1S).
+For all other frequencies (5.5, 9.0, 33.0, 35.0, 0.842, and 0.943 GHz), we report 5-sigma upper limits (see Table 8 of the appendix).
+The GRB position was also observed with the Australian SKA Pathfinder by the Rapid ASKAP Continuum Survey (RACS; 2020PASA...37...48M) at 0.842, and 0.943 GHz on 2023 July 1 and 2024 Jan 5.
+We find no radio detections in preliminary data and report 5-sigma upper limits of 1.5 mJy at 0.842 and 0.943 GHz.
+There were also no archival radio detection pre-burst within 1 arcminute in surveys: the National Radio Astronomy Observatory VLA Sky Survey (1998AJ....115.1693C) at 1.4 GHz, the Sydney University Molonglo Sky Survey (2003MNRAS.342.1117M) at 843 MHz, the Very Large Array Sky Survey (2020PASP..132c5001L) at 3 GHz or the RACS survey.
 We measured a VLASS 5-sigma upper limit of 0.69 mJy at 3 GHz.
@@
 Swift/XRT did not observe the location of X-ray afterglow of GRB 230204B until 91.5 ks post-burst.
-Therefore to search for the hard X-ray early afterglow, we used the BatAnalysis python package (Parsotan et al., 2023) to analyze Swift BAT survey data from 2023-02-03 to 2023-02-11.
+Therefore, to search for the hard X-ray early afterglow, we used the BatAnalysis python package (BatAnalysis) to analyze Swift BAT survey data from 2023-02-03 to 2023-02-11.
 We analyzed survey data, where the location of the GRB had at least a partial coding fraction of on the BAT detector plane.
 As expected, the GRB afterglow was not detected in the BAT survey data, which is typical since hard X-ray afterglows have never been detected at such late times.
-However, we were able to place 5 upper limits on the emission in the 14-195 keV energy band (shown in Figure 2).
-We additionally conducted a mosaiced analysis of the BAT survey data in 1-day time bins, for two days after the trigger time, and were able to place 5 flux upper limits on the emission in the 14-195 keV energy band where the flux of the GRB afterglow is erg/sec/cm2.
+However, we were able to place 5 upper limits on the emission in the 14-195 keV energy band.
+In Figure 3, we show the flux upper limits derived from individual BAT survey snapshots, which correspond to shorter time bins and thus shallower limits.
+To obtain more stringent constraints, we also performed a mosaiced analysis using 1-day bins for the two days after the trigger, yielding deeper upper limits of erg sec-1 cm-2.
+Although the BAT survey provides the deepest available constraints in the 14–195 keV range during this period, the obtained limits are not sufficiently deep to impose meaningful physical constraints on the afterglow.
+In particular, the derived upper limits are significantly above the flux level expected from a simple extrapolation of the contemporaneous Swift/XRT spectrum to hard X-rays.
+Thus, while these upper limits confirm the absence of an unusually bright hard X-ray component, they are too shallow to provide additional constraints on the standard afterglow emission.
 Based on the above observations, we have shown the broadband composite afterglow light curve of GRB 230204B in Figure 3.
 3 Results
-3.1 Prompt emission temporal and spectral characterize
+3.1 Prompt emission temporal and spectral characterization
 For the analysis of Fermi GBM data, we utilized the Fermi GBM Data Tools222https://fermi.gsfc.nasa.gov/ssc/data/analysis/gbm/gbm_data_tools/gdt-docs/ to extract temporal information.
 We selected specific NaI (NaI 7 and NaI 8) and BGO (BGO 1) detectors for the analysis based on their optimal viewing angles relative to the GRB’s position to maximize the signal-to-noise ratio.
-For the NaI detectors, we focused on two energy bands: 8-30 keV and 50-300 keV, allowing us to explore both low and high-energy photon populations.
+For the NaI detectors, we focused on two energy bands: 8-50 keV and 50-300 keV, allowing us to explore both low and high-energy photon populations.
+The 8-50 keV band was chosen to provide a more comprehensive coverage of the soft X-ray emission, while the 50-300 keV band captures the higher-energy component.
 For BGO 1, two bands were used: 0.3-1 MeV and 1-40 MeV, to investigate higher-energy photons.
@@
 This GRB exhibits four distinct emission episodes, each separated by periods of quiescence, with the shaded regions highlighting the intervals selected for time-integrated spectral analysis.
+The time windows (pulse 1: T0- 0.80 to T0 + 4.97 sec, pulse 2: T0 + 37.00 to T0 + 68.15 sec, pulse 3: T0 + 109.45 to T0 + 161.90 sec, and pulse 4: T0 + 172.58 to T0 + 229.69 sec) for these intervals were determined using the Bayesian Block algorithm, which adaptively identifies significant temporal structures in the light curve based on statistical criteria.
 The energy-dependent light curves show that the temporal structure and intensity vary significantly across different energy bands, reinforcing the diverse nature of GRB light curves.
@@
 The inset plot shows the MAXI light curve, covering the 2-20 keV energy range, further subdivided into 2-4 keV, 4-10 keV, and 10-20 keV.
-Although MAXI detected only part of the GRB, its observations provide valuable insights into the lower-energy emission and suggest a very softer nature of the last pulse of GRB 230204B.
+Although MAXI detected only part of the GRB, its observations provide valuable insights into the lower-energy emission and suggest a much softer nature of the last pulse of GRB 230204B.
 We analyzed the time-integrated (T0-0.80 to T0 +229.69 sec) Fermi spectrum of GRB 230204B using the Multi-Mission Maximum Likelihood framework (3ML) to evaluate its spectral characteristics.
-We explore several phenomenological models, including the Band function, Cutoff power-law (CPL), and combinations with a Blackbody component, which is theoretically expected to originate from the photospheric emission of GRBs.
+We explore several phenomenological models, including the Band function, Cutoff power-law (CPL), and combinations with a Blackbody component, which is theoretically expected to originate from the photospheric emission of GRBs (2010ApJ...709L.172R; 2013ApJ...765..103L).
 For each model, we employed Bayesian spectral fitting to obtain the Deviance Information Criterion (DIC) values, allowing us to systematically compare the model fits.
@@
 | 37.00-68.15 | 5083.02 | 5091.67 | 5058.11 | 5077.71 |
-| 109.450-161.90 | 5655.64 | 5660.37 | 5619.57 | 5606.26 |
-| 172.58-229.69 | 5601.18 | 5606.86 | 5575.45 | 5577.01 |
+| 109.45-161.90 | 5655.64 | 5660.37 | 5620.64 | 5606.26 |
+| 172.58-229.69 | 5601.18 | 5606.86 | 5577.01 | 5575.45 |
 | Time-intervals | Best-fit parameters | |||||
@@
 GRBs are known for their significant spectral evolution and high variability, with emission properties often varying notably across individual pulses within multi-pulse bursts.
-Li et al.
-(2021) conduct a comprehensive time-resolved Bayesian spectral analysis of well-defined GRB pulses observed by Fermi.
+Liang2021 conduct a comprehensive time-resolved Bayesian spectral analysis of well-defined GRB pulses observed by Fermi.
 They observe a trend toward spectral softening over time, largely reflected in the gradual decrease of the low-energy power-law index.
-We compared the time-resolved spectral analysis of GRB 230204B with the sample studied by Li et al.
-(2021).
+We compared the time-resolved spectral analysis of GRB 230204B with the sample studied by Liang2021.
 We also observe that the low-energy power-law indices of GRB 230204B become softer over time (hard to soft), with a distinct spectrum evolution from the first to later episodes (see Figure 5).
-Such spectral evolution feature of prompt emission could be explained as photospheric emission is most commonly observed near the trigger time, while synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Li et al., 2021).
+Such a spectral evolution feature of prompt emission could be explained as photospheric emission is most commonly observed near the trigger time, while synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Liang2021).
 3.2 Pulse-wise prompt emission correlation and analysis
-The pulse-wise analysis of GRB prompt emission provides a valuable perspective on the Amati and Yonetoku correlations (key relationships used to explore the physical properties of GRBs), enhancing our understanding of GRB spectral evolution and energetics on a finer timescale (Chand et al., 2020).
+The pulse-wise analysis of GRB prompt emission provides a valuable perspective on the Amati and Yonetoku correlations (key relationships used to explore the physical properties of GRBs), enhancing our understanding of GRB spectral evolution and energetics on a finer timescale (2020ApJ...898...42C).
 By examining each pulse independently, rather than averaging over the entire prompt emission phase, we gain insights into the distinct characteristics of individual emission episodes within a single burst.
-In the Amati correlation, which relates the isotropic energy () to the spectral peak energy in the rest frame (Amati, 2006), each pulse may fall along or diverge from the established long and short GRB trends, providing information about the pulse-specific energy release mechanisms.
-Such an analysis not only clarifies the evolutionary trajectory of the GRB emission but also tests the applicability of the Amati correlation at sub-burst levels (Basak & Rao, 2013).
+In the Amati correlation, which relates the isotropic energy () to the spectral peak energy in the rest frame (2006MNRAS.372..233A), each pulse may fall along or diverge from the established long and short GRB trends, providing information about the pulse-specific energy release mechanisms.
+Such an analysis not only clarifies the evolutionary trajectory of the GRB emission but also tests the applicability of the Amati correlation at sub-burst levels (2013MNRAS.436.3082B).
 In the top-left panel of Figure 11, we illustrate the positioning of each pulse (plotted as colored squares) of GRB 230204B within the established Amati correlation.
-For reference, samples of long and short bursts from Minaev & Pozanenko (2020) are represented by orange and blue circles, respectively.
+For reference, samples of long and short bursts from 2020MNRAS.492.1919M are represented by orange and blue circles, respectively.
 The linear fits corresponding to these groups are indicated by solid orange and blue lines, with 3 uncertainties depicted by shaded bands around each fit.
-The time-integrated and pulse-wise analysis of GRB 230204B follows the Amati plane of long GRB populations.
-The top-right panel of Figure 10 presents GRB 230204B positions within the Yonetoku correlation framework, a relation between peak energy and the isotropic peak luminosity () in the rest frame (Yonetoku et al., 2004).
-Here, the short and long GRB samples from Nava et al.
-(2012) are shown by blue and orange circles, with the 3 scatter around each population’s best-fit line visualized by shaded areas.
-GRB 230204B follows the Yonetoku correlation and lies on the upper side of this correlation.
-The results in both correlations offer an insightful examination of GRB 230204B, highlighting the pulse-specific properties within established GRB behavior and indicating the energetic and luminous nature of the burst.
+The time-integrated ( = 1.92 erg) shows that GRB 230204B follows the Amati plane of long GRB populations.
+However, when considering the episodes individually, we find that while later pulse Episode 4 remains broadly consistent with the long GRB population, Episodes 1, 2, and 3 fall into an intermediate region between the long and short GRB distributions, indicating a less well-defined classification.
+The top-right panel of Figure 11 presents GRB 230204B positions within the Yonetoku correlation framework, a relation between peak energy and the isotropic peak luminosity () in the rest frame (2004ApJ...609..935Y).
+Here, the short and long GRB samples from 2012MNRAS.421.1256N are shown by blue and orange circles, with the 3 scatter around the population’s best-fit line visualized by a shaded area.
+Similar to the Amati plane, the overall burst lies on the upper side of the Yonetoku correlation for long GRBs.
 3.2.1 Spectral Peak-Duration Distribution Analysis
-Long-duration GRBs are generally characterized by a softer spectrum relative to the typically harder spectrum observed in the majority of short-duration GRBs, as demonstrated in previous studies (e.g., Kouveliotou et al.
-1993).
-To analyze this characteristic for GRB 230204B, we measured both the time-integrated and episode-wise spectral peak energy (an indicator of spectral hardness) alongside the duration (the time during which 90% of the burst’s total fluence is detected, for example, calculation for the second pulse of GRB 230204B is shown in the bottom left of Figure A1 of the appendix).
-We then plotted the pulse-wise values on the - plane to compare with a larger dataset of GRBs detected by the Fermi GBM, as illustrated in the bottom right panel of Figure 10.
-Our results indicate that the time-integrated and episode-wise properties of GRB 230204B exhibit softer values and durations exceeding 2 seconds, placing it on the right side of the spectral peak-energy versus duration plane.
-This position aligns well with the properties typically observed for long-duration GRBs, as they occupy the region in the - plane corresponding to longer durations and softer spectra.
-This comparison confirms that GRB 230204B follows the standard characteristics of long-duration GRBs.
+Long-duration GRBs are generally characterized by a softer spectrum relative to the typically harder spectrum observed in the majority of short-duration GRBs, as demonstrated in previous studies (e.g., 1993ApJ...413L.101K).
+To analyze this characteristic for GRB 230204B, we measured both the time-integrated and episode-wise spectral peak energy (an indicator of spectral hardness) alongside the duration (the time during which 90% of the burst’s total fluence is detected, values are 214.02 sec, 4.86 sec, 24.32 sec, 41.98 sec, and 51.46 sec for time-integrated duration, episode 1, episode 2, episode 3, and episode 4, respectively).
+We then plotted the pulse-wise values on the - plane to compare with a larger dataset of GRBs detected by the Fermi GBM, as illustrated in the bottom panel of Figure 11.
+Our results show that the time-integrated properties of GRB 230204B exhibit a relatively soft value and duration longer than 2 sec, placing the burst clearly within the Type II (collapsar) GRB population.
+However, the episode-wise analysis reveals interesting diversity in spectral and temporal properties: Episode 1 lies an intermediate region in both the - and Amati correlation (see Section 3.2) plane, with relatively short duration ( 4.86 sec), while Episodes 2 and 3 show progressive evolution toward longer durations and softer spectra.
+Episode 4 firmly places itself within the long GRB distribution with the softest spectrum ( -0.80) and longest duration (T90 51.46 sec).
+This diversity likely reflects evolving emission mechanisms throughout the burst.
+The spectral evolution from photosphere-dominated to synchrotron-dominated emission supports hybrid jet models where thermal and non-thermal components coexist (Liang2021).
+Alternatively, for collapsar with residual stellar envelopes, the jet’s interaction with material at different radii can produce distinct emission episodes with varying durations and spectral properties (2003ApJ...584..390W; 2003ApJ...586..356Z).
+Our analysis demonstrates that individual emission episodes within a single long GRB can exhibit durations and spectral properties overlapping with the short GRB population and provides evidence for complex, evolving emission physics throughout the burst duration.
+This finding adds to growing evidence that the traditional 2-second duration threshold for GRB classification may be overly simplistic and that progenitor identification requires consideration of additional properties beyond alone (2025arXiv250800142G).
+GRB 230204B exemplifies how such multi-layered approaches reveal complex emission physics that would remain hidden in time-integrated analyses alone.
+This underscores the importance of pulse-resolved studies for understanding the diversity of GRB emission mechanisms and jet physics.
 3.3 Afterglow analysis and correlation
-Based on the multi-wavelength follow-up observations campaign ( 1.3 ks to 335 days post burst) of GRB 230204B, we noted that BOOTES robotic and DOT observations have a well-sampled light curve in optical / filters.
-We fitted flux density optical /-band light curves using a simple power-law function with a temporal decay index of -1.85 0.03 (/DOF = 43.47/52).
-Furthermore, we combined all the optical data (MASTER, BOOTES, DOT, and a few points from GCN) to get a more well-sampled light curve and it could also be fitted with a simple power-law function with a temporal decay index of -1.84 0.02 (/DOF = 81.13/58).
-We noted that the optical temporal index is steeper than the typically observed decay slope of -1 (Dainotti et al., 2022).
-On the other hand, the X-ray afterglow light curve of GRB 230204B has very limited (close to the sensitivity of XRT) data points as Swift followed the burst during the late phase and could be fitted with a temporal decay index of -0.72 0.63 (/DOF = 0.58/1).
+Based on the multi-wavelength follow-up observations campaign ( 1.3 ks to 335 days post burst) of GRB 230204B, we noted that BOOTES robotic and DOT observations have a well-sampled light curve in optical and filters, respectively.
+We fitted flux density optical and r-bands light curves using a simple power-law function with a temporal decay index of -1.67 0.15 (/DOF = 20.17/47) and -1.83 0.05 (/DOF = 8.25/2), respectively.
+Furthermore, we combined all the optical data (MASTER, BOOTES, DOT, and a few points from GCN) to get a more well-sampled light curve, and it could also be fitted with a simple power-law function with a temporal decay index of -1.84 0.02 (/DOF = 81.13/58).
+We noted that the optical temporal index is steeper than the typically observed decay slope of -1 (2022ApJ...940..169D).
+On the other hand, the X-ray afterglow light curve of GRB 230204B has very limited (the last XRT data point has a marginal detection with 3.53) data points as Swift followed the burst during the late phase and could be fitted with a temporal decay index of -0.72 0.63 (/DOF = 0.58/1).
 The correlation between the average optical decay index and optical luminosity provides crucial insights into the physical mechanisms governing GRB afterglows and their surrounding environments.
-Oates et al.
-(2012, 2015) examined the possible correlation between the average optical temporal decay indices and optical luminosity (at 200-sec post burst) for GRBs with optical afterglows observed by Swift/UVOT.
-They found evidence of anti-correlation between these two parameters (200-sec post burst) as shown in Figure 5.
+2012MNRAS.426L..86O; 2015MNRAS.453.4121O examined the possible correlation between the average optical temporal decay indices and optical luminosity (at 200-sec post-burst) for GRBs with optical afterglows observed by Swift/UVOT.
+They found evidence of anti-correlation between these two parameters (200-sec post-burst) as shown in Figure 6.
 This correlation indicates that optical luminous bursts decay more rapidly than optical faint bursts.
 We measured optical luminosity and temporal decay index for GRB 230204B and overlaid it onto the luminosity-decay correlation.
-GRB 230204B exhibits a relatively steep decay rate ( 1.4 in MASTER filter) and high optical luminosity (at the first epoch of afterglow detection).
+GRB 230204B exhibits a relatively steep decay rate ( 1.74 average decay index) and high optical luminosity (at the first epoch of afterglow detection).
 We noted GRB 230204B, highlighted in red in the figure, aligns well with the global correlation observed for long-duration GRBs and lies towards the edge (see Figure 6).
@@
 3.4 Broadband afterglow modeling of GRB 230204B
-| Model | Top hat | Gaussian | Gaussian Core |
-| 0.003 | 0.004 | 0.005 | |
-| 0.006 | 0.006 | 0.007 | |
-| – | 4.020 | 4.714 | |
-| 55.621 | 55.529 | 55.143 | |
-| 2.280 | 2.267 | 2.252 | |
-| 1.449 | 1.987 | 2.685 | |
-| -5.669 | -5.899 | -6.133 | |
-| -0.525 | -0.353 | -0.198 | |
-| 1.0 | 1.0 | 1.0 | |
-| ln () | -101.310.08 | -107.500.08 | -110.530.08 |
-| ln (BF) | 7638.16 | 7631.97 | 7628.94 |
+| Model | Prior | Top hat | Gaussian | Gaussian Core |
+| [0, /2] | 0.003 | 0.004 | 0.005 | |
+| [0.0001, 0.1] | 0.006 | 0.006 | 0.007 | |
+| [1, 8] | – | 4.020 | 4.714 | |
+| [52, 57] | 55.621 | 55.529 | 55.143 | |
+| [2.0, 3.5] | 2.280 | 2.267 | 2.252 | |
+| ) | [-3, 4] | 1.449 | 1.987 | 2.685 |
+| [-7, 0] | -5.669 | -5.899 | -6.133 | |
+| [-5, 0] | -0.525 | -0.353 | -0.198 | |
+| 1.0 | 1.0 | 1.0 | ||
+| ln | -101.310.08 | -107.500.08 | -110.530.08 | |
+| ln BF | 7638.16 | 7631.97 | 7628.94 |
 The broadband afterglow modeling involved fitting the observed data across multiple wavelengths, including radio, optical, and X-ray bands.
 Such broadband modeling helps to understand the physical parameters of the jet, such as its opening angle, energy, and the surrounding medium’s density.
-The afterglow emission is generally explained by the synchrotron radiation model (Sari et al., 1998; Meszaros & Rees, 1997b; Sari & Piran, 1999).
+The afterglow emission is generally explained by the synchrotron radiation model (1998ApJ...497L..17S; mes97; sari99).
 In this model, relativistic electrons are accelerated by the GRB shockwave spiral in the magnetic field, emitting radiation across the electromagnetic spectrum.
 The synchrotron spectrum is characterized by several break frequencies (self-absorption, peak, and cooling frequencies), which evolve with time and affect the observed light curves at different wavelengths.
-For the afterglow modeling of GRB 230204B, we utilized afterglowpy333It currently does not account for certain features, including reverse shock emission, an external wind medium (n r-2), or synchrotron self-absorption.
-(Ryan et al., 2020)/Redback (Sarin et al., 2024) to explore various jet structure models, including the Tophat, Gaussian, and GaussianCore models.
+For the afterglow modeling of GRB 230204B, we performed fitting using Redback software (2024MNRAS.531.1203S) via importing afterglow jet models from afterglowpy333It currently does not account for certain features, including reverse shock emission, an external wind medium (n r-2), or synchrotron self-absorption.
+(2020ApJ...896..166R) to explore various jet structure models, including the Tophat, Gaussian, and GaussianCore models.
 The Bayesian inference framework provided by Redback allowed us to test different priors and achieve robust parameter estimation.
 Specifically, we employed the PyMultiNest sampler to explore the parameter space and used well-defined priors for key parameters such as the observer angle (), core angle (), initial energy (E0), and interstellar medium density (nISM).
-Due to the limited number of data points, we have fixed bulk Lorentz factor ( 800), constrained using the -Lorentz factor correlation (Liang et al., 2010).
+Due to the limited number of data points, we have fixed the bulk Lorentz factor ( 800), constrained using the -Lorentz factor correlation (2010ApJ...725.2209L).
 The priors used in the analysis include uniform distributions for the energy, core angle, ambient medium density, and a sine prior for the observer’s viewing angle.
@@
 The inferred initial energy is relatively high at 55.62 erg, and the ISM density is constrained to 1.45 cm-3.
+We also modeled the afterglow of GRB 230204B with VegasAfterglow (2025arXiv250710829W), assuming a top-hat jet in an ISM-like medium, and obtained parameters consistent with those derived using afterglowpy.
+The best-fit values are: rad, rad, erg, , , , and .
 The Top-Hat jet model provides a robust description of GRB 230204B’s afterglow, indicating a narrow, highly collimated jet.
@@
 The tightly constrained parameters for and indicate that we are likely observing the jet from a viewing angle close to the core, providing a relatively straightforward interpretation in the context of synchrotron emission from a top-hat jet model.
+The last XRT data point at 10 days post-burst shows a deviation from the afterglow model prediction.
+This observation corresponds to a marginal detection (3.53) with a count rate of ct sec-1 (exposure of 4533.2 sec), close to the sensitivity limit of Swift-XRT.
+Due to the limited photon statistics, we could not reliably constrain the spectral parameters.
+Considering the excess flux is intrinsic, it could arise from additional physical processes such as late-time energy injection (2006ApJ...647.1238J), refreshed shocks (2022MNRAS.513.2777K), or wavelength-dependent emission components.
+However, in the absence of further X-ray or multiwavelength coverage at comparable epochs, the true origin of this excess remains uncertain.
 A notable feature in the fit is the very low value of , the magnetic field fraction, which was constrained to = -5.67.
-This low magnetic field can plausibly arise from shock compression of the ambient interstellar medium (ISM), which supports a weakly magnetized medium (Kumar & Barniol Duran, 2009; Wang et al., 2015).
-Similar values of have been observed in GRBs such as GRB 190530A (Gupta et al., 2022a), GRB 171205A (Leung et al., 2021), GRB 130427A (Fraija et al., 2016), and a few others (Kumar & Barniol Duran, 2009, 2010; Santana et al., 2014), highlighting the diversity in GRB environments.
-For GRBs, when , there might be Synchrotron Self-Compton (SSC) contribution to afterglow, particularly at higher frequencies (Joshi & Razzaque, 2021).
-However, in the case of GRB 230204B, the SSC effects appear negligible within the observed wavelength range (i.e., low-frequencies radio and optical), consistent with findings for similar GRBs (Gupta et al., 2022a; Wang et al., 2024).
-This is further supported by the temporal decay slopes, afterglow modeling (afterglowpy primarily models synchrotron emission processes), and lack of GeV emission (consistent with synchrotron-dominated regimes).
+Such low magnetic fields can possibly arise from shock compression of a weakly magnetized ambient medium (2009MNRAS.400L..75K; 2015ApJS..219....9W), and similar values have been observed in GRBs such as GRB 190530A (Gupta1), GRB 171205A (2021MNRAS.503.1847L), GRB 130427A (2016ApJ...818..190F) and a few others (2009MNRAS.400L..75K; 2010MNRAS.409..226K; 2014ApJ...785...29S), highlighting the diversity in GRB environments.
+For GRBs, when , there might be Synchrotron Self-Compton (SSC) contribution to afterglow (2021MNRAS.505.1718J).
+However, in the case of GRB 230204B, our afterglow modeling using Redback and VegasAfterglow (see below) suggests that the SSC effects appear negligible, consistent with findings for similar GRBs (Gupta1; 2024ApJ...969..146W).
+This is further supported by the temporal decay slopes, and no significant GeV emission (consistent with synchrotron-dominated regimes).
+However, we cannot completely rule out some SSC contribution, particularly at early radio epochs, where current limited data may lack the sensitivity to detect weak SSC components.
 Future studies focusing on a broader sample of bright GRBs with low , combined with well-sampled broadband light curves (from gamma-ray to radio) and afterglow modeling that includes SSC contributions, have the potential to shed light on their physical origins and diverse emission mechanisms.
+Furthermore, to investigate alternative scenarios, we modeled the afterglow of GRB 230204B using the VegasAfterglow framework (2025arXiv250710829W), a high-performance and versatile code that self-consistently treats forward and reverse shock dynamics, synchrotron and inverse Compton radiation, and arbitrary circumburst density profiles.
+Unlike commonly used tools such as afterglowpy, VegasAfterglow supports wind-like media, allowing us to probe environments shaped by massive-star mass loss.
+We performed Markov Chain Monte Carlo (MCMC) fitting of the broadband afterglow light curves under the assumption of a stellar-wind environment (see Table 4) with several jet structures (top-hat, power-law, and Gaussian).
+The best-fitting wind model with a top-hat jet and resulting posterior distributions for key physical parameters are shown in Figures 13 and 14 of the appendix, respectively.
+This modeling yields constraints on the isotropic-equivalent energy, jet opening angle, electron distribution index, and other microphysical parameters, indicating a luminous afterglow produced by a narrowly collimated jet.
+To discriminate between ISM and wind-like circumburst environments for GRB 230204B, we performed a statistical comparison using reduced values for both the models.
+For the ISM top-hat model: , and for the Wind top-hat model: .
+While both reduced values are greater than 1, the ISM model provides a notably better fit.
+To quantify the statistical significance of this improvement, we applied an F-test.
+The F-statistic yields (F-value), and corresponds to a p-value of , indicating that the ISM model is statistically significantly better at the confidence level.
+This indicates that the improvement is not merely due to random fluctuations but reflects a genuine preference for the ISM scenario.
+Beyond the statistical framework, we examined the standard closure relations for model discrimination.
+Using the electron energy distribution indices obtained from our afterglow modeling: (ISM scenario), (wind scenario), we calculated the predicted temporal and spectral indices.
+For the slow-cooling relations, the closure relations are:
+For the ISM-derived , these relations predict or and temporal slopes or for the () and () spectral regimes, respectively.
+For the wind-derived , we obtain or , and temporal indices or for the analogous spectral regimes.
+Notably, both ISM and wind predictions systematically underestimate the observed steep optical decay rate.
+Consequently, the standard closure relations fail to discriminate the environments, primarily due to the unavailability of well-sampled multiwavelength data.
+In contrast, our comprehensive afterglow modeling framework provides a more robust and physically complete determination that the ISM scenario is preferred.
 4 Discussion
 4.1 Spectral evolution and Possible Radiation Mechanism
-The spectral evolution of the prompt emission reveals complex temporal and spectral variability, which provides crucial insights into the nature of the emission mechanisms and the physical conditions in the GRB jet (Bošnjak & Daigne, 2014).
+The spectral evolution of the prompt emission reveals complex temporal and spectral variability, which provides crucial insights into the nature of the emission mechanisms and the physical conditions in the GRB jet (2014A&A...568A..45B).
 Typically, GRB prompt spectra evolve over time, with the peak energy often shifting to lower energies as the burst progresses.
-This trend, known as “hard-to-soft” evolution, is commonly observed and is thought to reflect cooling processes within the jet (Norris et al., 1986; Bhat et al., 1994).
-Alternatively, “tracking” behavior, where follows the intensity variations of the burst, has also been observed in some cases (Ryde & Svensson, 1999), suggesting a possible relationship between the central engine activity and the observed emission (Gupta et al., 2021b).
+This trend, known as “hard-to-soft” evolution, is commonly observed and is thought to reflect cooling processes within the jet (1986ApJ...301..213N; 1994ApJ...426..604B).
+Alternatively, “tracking” behavior, where follows the intensity variations of the burst, has also been observed in some cases (1999ApJ...512..693R), suggesting a possible relationship between the central engine activity and the observed emission (2021MNRAS.505.4086G).
 The radiation mechanisms driving GRBs are fundamentally linked to the spectral index observed during the prompt emission phase and the overall radiative efficiency.
 The spectral index , which describes the shape of the GRB’s spectrum, provides crucial insights into the underlying physical processes.
-A harder low energy spectral index (i.e., closer to or exceeding -2/3) typically suggests a dominant photospheric or thermal component, where radiation emerges directly from the photosphere of the relativistic outflow.
-In contrast, a softer a (less than -2/3) often points to non-thermal processes, such as synchrotron emission, where electrons are accelerated in magnetic fields, producing a broad power-law spectrum (Chattopadhyay et al., 2022; Caballero-García et al., 2023).
+A harder low-energy spectral index (i.e., closer to or exceeding -2/3) typically suggests a dominant photospheric or thermal component, where radiation emerges directly from the photosphere of the relativistic outflow.
+In contrast, a softer (less than -2/3) often points to non-thermal processes, such as synchrotron emission, where electrons are accelerated in magnetic fields, producing a broad power-law spectrum (2022ApJ...936...12C; 2023MNRAS.519.3201C).
 To investigate the spectral evolution of GRB 230204B, we conducted a time-resolved spectral analysis, selecting time bins based on a Bayesian algorithm for optimal selection.
@@
 These parameters are overlaid on the count-rate light curve of the burst’s prompt emission, revealing that the observed exhibits an intensity-tracking behavior.
-Harder spectral indices during the first and second episodes might suggest a prominent photospheric component, which often produces harder spectra due to thermal-like emission (Ryde et al., 2010).
-The softer spectral indices during the later bins of the second, third, and fourth episodes might suggest a synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Li et al., 2021; Acuner & Ryde, 2018).
-Additionally, we explored correlations among the spectral parameters, examining (a) log(Flux) versus log(), (b) log(Flux) versus , and (c) log() versus .
-Pearson correlation analysis was used to quantify the correlation strengths (Pearson correlation coefficient, (r) and to assess the statistical significance (p-value) for each parameter pair.
+Harder spectral indices during the first and second episodes might suggest a prominent photospheric component, which often produces harder spectra due to thermal-like emission (2010ApJ...709L.172R).
+The softer spectral indices during the later bins of the second, third, and fourth episodes might suggest that a synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Liang2021; 2018MNRAS.475.1708A).
+Additionally, we explored correlations among the spectral parameters, examining 1.
+log(Flux) versus log(), 2.
+log(Flux) versus , and 3.
+log() versus .
+Pearson correlation analysis was used to quantify the correlation strengths (Pearson correlation coefficient, r) and to assess the statistical significance (p-value) for each parameter pair.
 We noted a strong correlation for log(Flux) versus log() of Band and CPL models with correlation coefficients equal to 0.69 and 0.73, respectively.
@@
 The radiation efficiency of GRBs is another key factor, reflecting the fraction of the jet’s kinetic energy converted into gamma rays.
-It varies significantly depending on the dominant radiation mechanisms within the jet (Wang et al., 2015).
+It varies significantly depending on the dominant radiation mechanisms within the jet (2015ApJS..219....9W).
 In the photospheric model, radiation is produced as photons escape from the optically thick inner regions of the relativistic outflow.
-Energy is released at the photosphere where the flow becomes transparent, leading to thermal emission.
-Photospheric models tend to predict higher radiation efficiencies, sometimes exceeding 50-60%, especially if energy dissipation occurs near the photosphere (Lazzati et al., 2009; Beloborodov, 2010; Pe’er & Ryde, 2011; Lazzati et al., 2013).
-In contrast, the synchrotron models (arising from accelerated particles in internal shocks or magnetic reconnection sites) often predict lower efficiencies, typically in the range of a few percent to–30%, in such cases much of the energy remains in the kinetic form rather than being radiated away (Kobayashi et al., 1997; Spada et al., 2000; Beniamini & Piran, 2013; Kumar & Zhang, 2015).
+Energy is released at the photosphere, where the flow becomes transparent, leading to thermal emission.
+Photospheric models tend to predict higher radiation efficiencies, sometimes exceeding 50-60%, especially if energy dissipation occurs near the photosphere (2009ApJ...700L..47L; 2010MNRAS.407.1033B; 2011ApJ...732...49P; 2013ApJ...765..103L).
+In contrast, the synchrotron models (arising from accelerated particles in internal shocks or magnetic reconnection sites) often predict lower efficiencies, typically in the range of a few percent to 30%, in such cases much of the energy remains in the kinetic form rather than being radiated away (1997ApJ...490...92K; 2000ApJ...537..824S; 2013ApJ...769...69B; 2015PhR...561....1K).
 The hybrid model combines thermal photospheric emission with non-thermal synchrotron radiation from shocks or magnetic reconnections within the jet.
 This model can accommodate both thermal and non-thermal features in GRB spectra.
-Hybrid models can achieve efficiencies ranging from a few percent to 50% or higher, as they leverage both thermal and non-thermal emission processes (Zhang & Yan, 2011; Zhang & Pe’er, 2009; Vurm et al., 2011).
+Hybrid models can achieve efficiencies ranging from a few percent to 50% or higher, as they leverage both thermal and non-thermal emission processes (2011ApJ...726...90Z; 2009ApJ...700L..65Z; 2011ApJ...738...77V).
 These discussions suggest that GRBs may involve a complex interplay of thermal and non-thermal processes, with the relative contributions of these mechanisms varying across different bursts and even between different phases of the same burst.
@@
 Our analysis of the prompt emission reveals a spectral hardening during the initial pulses, with a subsequent softening in later pulses, consistent with a hybrid composition (baryonic + Poynting-flux) of Band + Blackbody components.
-This spectral evolution suggests that the jet may consist of both photospheric (thermal) and synchrotron (non-thermal) components, with the harder spectrum at the start dominated by a photospheric component and later phases increasingly influenced by synchrotron radiation (Sharma et al., 2020; Gupta et al., 2024c).
+This spectral evolution suggests that the jet may consist of both photospheric (thermal) and synchrotron (non-thermal) components, with the harder spectrum at the start dominated by a photospheric component and later phases increasingly influenced by synchrotron radiation (2020MNRAS.493.5218S; 2024ApJ...972..166G).
 Additionally, our afterglow modeling indicates a low gamma-ray efficiency of 4.3%, suggesting that most of the jet’s energy remains in kinetic form with minimal conversion into prompt gamma-ray emission.
-While the low gamma-ray efficiency supports a hybrid model, where both thermal photospheric and non-thermal synchrotron components coexist, it does not entirely rule out a synchrotron-only scenario (Rossi et al., 2002).
+While the low gamma-ray efficiency supports a hybrid model, where both thermal photospheric and non-thermal synchrotron components coexist, it does not entirely rule out a synchrotron-only scenario (2002MNRAS.332..945R).
 Such a model, without any photospheric thermal emission, could also account for the low efficiency, provided the energy dissipation mechanism predominantly drives non-thermal radiation.
@@
 Furthermore, the low gamma-ray efficiency aligns well with the characteristics of an ISM-like top-hat jet structure in the afterglow phase.
-In such a scenario, a top-hat jet interacting with a homogeneous interstellar medium produces a relatively uniform energy distribution across the jet, reducing the likelihood of highly efficient gamma-ray emission (Sari et al., 1998; Granot & Sari, 2002; Zhang & Kobayashi, 2005).
+In such a scenario, a top-hat jet interacting with a homogeneous interstellar medium produces a relatively uniform energy distribution across the jet, reducing the likelihood of highly efficient gamma-ray emission (1998ApJ...497L..17S; 2002ApJ...568..820G; 2005ApJ...628..315Z).
 This jet structure also favors an afterglow powered predominantly by forward shock emission, which is consistent with the observed light curve and its steady decay pattern.
@@
 The kinetic energy for GRB 230204B, inferred from afterglow modeling, was found to be significantly higher than the isotropic gamma-ray energy, emphasizing that a substantial fraction of the jet’s energy remained in kinetic form and was not radiated as gamma rays.
-We noted that GRB 230204B has a higher and values than expected from magnetar central engine (2 1052 erg, Metzger et al.
-2011).
-This supports the rapidly rotating black hole as a possible central engine for this energetic and bright burst (Li et al., 2018; Ror et al., 2024).
+We noted that GRB 230204B has a higher and values than expected from magnetar central engine (2 1052 erg, 2011MNRAS.413.2031M).
+This supports the rapidly rotating black hole as a possible central engine for this energetic and bright burst (2018ApJS..236...26L; 2024ApJ...971..163R).
 The distribution of radiation efficiencies across the GRB population, as shown in Figure 9, highlights GRB 230204B as the upper edge with a very low efficiency.
 4.2 Brightness Comparison to GRB sample and physical scenarios for GRB 230204B
-The optical luminosity of GRB 230204B (R/r-band and g-band data) has been compared with the population of GRBs with measured redshifts (Dainotti et al., 2024), providing valuable insights into its optical emission characteristics.
+In this Section, we compared the optical, X-ray, and radio afterglow brightness of GRB 230204B with a larger population of GRBs with measured redshifts and explored the physical scenarios.
+The optical luminosity of GRB 230204B (R/r-band and g-band data) has been compared with the population of GRBs with measured redshifts (2024MNRAS.533.4023D), providing valuable insights into its optical emission characteristics.
 The temporal evolution of GRB 230204B’s optical luminosity (see Figure 10) reveals an exceptionally bright initial emission, followed by a decay phase consistent with a forward shock model in a homogeneous interstellar medium (ISM).
 This places it among the brightest optical GRBs ever observed at the epoch of its detection, making it a standout event in the context of optical afterglow studies.
-The extreme optical brightness is an order of magnitude higher than other highly energetic bursts ( 1.3 ks post-burst) such as GRB 080319B (Racusin et al., 2008; Pandey et al., 2009), GRB 990123 (Castro-Tirado et al., 1999), GRB 190114C (Misra et al., 2021; Gupta et al., 2021a), and GRB 221009A.
-Furthermore, we have also examined the optical luminosity of GRB 230204B at the epoch of its detection as a function of redshift for a population of GRBs with measured redshifts (Dainotti et al., 2024).
+The extreme optical brightness is an order of magnitude higher than other highly energetic bursts ( 1.3 ks post-burst) such as GRB 080319B (2008Natur.455..183R; 2009A&A...504...45P), GRB 990123 (1999Sci...283.2069C), GRB 190114C (2021MNRAS.504.5685M; 2021RMxAC..53..113G), and GRB 221009A.
+Furthermore, we have also examined the optical luminosity of GRB 230204B at the epoch of its detection as a function of redshift for a population of GRBs with measured redshifts (2024MNRAS.533.4023D).
 We noted that GRB 230204B has the highest optical luminosity ( 1.1-2.4 ks post-burst) with respect to all known GRBs with measured redshifts at the same epoch (see Figure 10).
 This suggests that GRB 230204B could represent a highly efficient energy transfer to the optical afterglow, or it might be related to the properties of its surrounding environment, jet structure, or progenitor system, contributing to its extraordinary luminosity.
-Additionally, we noted that GRB 230204B follows an anti-correlation between the optical decay index and optical luminosity (Oates et al., 2012, 2015).
+Additionally, we noted that GRB 230204B follows an anti-correlation between the optical decay index and optical luminosity (2012MNRAS.426L..86O; 2015MNRAS.453.4121O).
 There are several scenarios where such a correlation could be explained.
-Continued energy injection from a central engine, such as a magnetar or a rapidly spinning black hole, could initially produce a very bright afterglow (Zhang & Mészáros, 2001).
-However, if the injection is short-lived or if the engine rapidly loses power, the afterglow could exhibit a steep decay as the additional energy source wanes (Zhang & Mészáros, 2001; Dai & Lu, 1998).
+Continued energy injection from a central engine, such as a magnetar or a rapidly spinning black hole, could initially produce a very bright afterglow (2001ApJ...552L..35Z).
+However, if the injection is short-lived or if the engine rapidly loses power, the afterglow could exhibit a steep decay as the additional energy source wanes (2001ApJ...552L..35Z; 1998A&A...333L..87D).
 Our afterglow analysis did not show evidence of central engine activity (flare, plateau, variability, etc).
-A reverse shock could also produce a very luminous optical flash (Zhang & Kobayashi, 2005).
+A reverse shock could also produce a very luminous optical flash (2005ApJ...628..315Z).
 If the reverse shock is strong, it can dominate the early afterglow.
@@
 For GRB 230204B, however, our afterglow modeling suggests that a forward shock model could explain the early optical emission.
-Another possibility involves a structured jet, where a bright core with steep wings might cause a steep decay if the emission transitions from being core-dominated to wing-dominated, or if there is a slight change in the viewing angle (Rossi et al., 2002; Granot & Kumar, 2003).
+Another possibility involves a structured jet, where a bright core with steep wings might cause a steep decay if the emission transitions from being core-dominated to wing-dominated, or if there is a slight change in the viewing angle (2002MNRAS.332..945R; 2003ApJ...591.1086G).
 However, our afterglow modeling suggests that a top-hat jet could explain the afterglow of GRB 230204B, so we discarded the structured jet for this burst.
-If the GRB is in a wind-like medium (density r-2), the early afterglow can be very bright (Chevalier & Li, 2000).
-A change in the density profile (e.g., transition to a uniform medium) could cause the afterglow to decay steeply (Panaitescu & Kumar, 2002; Starling et al., 2008).
-However, for GRB 230204B, our afterglow modeling suggests that an ISM-like medium could explain the broadband emission.
-A jet break occurs when the relativistic beaming angle of the jet becomes wider than the observer’s line of sight (Racusin et al., 2009).
-This typically results in a steepening of the afterglow light curve (Liang et al., 2008).
-An achromatic temporal slope of -2 at late times could indicate that the break occurred, however, for GRB 230204B, synchronous broadband data is not available to confirm such behavior, also such an early jet break is not expected so we discarded the possibility.
-GRB afterglows depend strongly on the observer’s viewing angle relative to the jet’s core (Granot et al., 1999; van Eerten & MacFadyen, 2013).
+If the GRB is in a wind-like medium (density r-2), the early afterglow can be very bright (2000ApJ...536..195C).
+A change in the density profile (e.g., a transition from a wind to a uniform medium) can also produce a steep decay in the afterglow light curve (2002ApJ...571..779P; 2008ApJ...672..433S).
+For GRB 230204B, we modeled the afterglow in both ISM and wind-like environments.
+While the wind scenario can reproduce some features of the light curve, our broadband analysis shows that the ISM-like medium provides a better overall fit to the data.
+A jet break occurs when the relativistic beaming angle of the jet becomes wider than the observer’s line of sight (2009ApJ...698...43R).
+This typically results in a steepening of the afterglow light curve (2008ApJ...675..528L).
+An achromatic temporal slope of -2 at late times could indicate that the break occurred; however, for GRB 230204B, synchronous broadband data is not available to confirm such behavior, and such an early jet break is not expected, so we discarded the possibility.
+GRB afterglows depend strongly on the observer’s viewing angle relative to the jet’s core (1999ApJ...513..679G; 2013ApJ...767..141V).
 When an observer views the GRB slightly off-axis but still close to the jet core, the observed afterglow appears very bright initially.
 This is due to the deceleration of the relativistic jet as it interacts with the surrounding medium, causing the beamed emission to broaden and spread into the observer’s line of sight.
-As the jet slows further and its emission becomes more isotropic, the light curve steepens, showing a rapid decline (Ryan et al., 2015).
-Such behavior suggests that GRB 230204B is observed close to the jet axis, as indicated by its narrow inferred viewing angle from afterglow modeling.
-The steep early optical decay and extreme luminosity support this interpretation.
-In Figure 12, we present a comparative analysis of the X-ray luminosity light curves of GRB 230204B with a sample of other GRBs that have measured redshifts (Gupta et al., 2024b).
+As the jet slows further and its emission becomes more isotropic, the light curve steepens, producing a rapid decline (2015ApJ...799....3R).
+For GRB 230204B, our afterglow modeling yields an observer angle of rad and a jet core angle of rad.
+Although the nominal values of the observer and jet core angles suggest that the line of sight is not exactly aligned with the jet axis.
+However, within the reported uncertainties, the two angles are consistent.
+This is possibly compatible with the steep early optical decay and high luminosity.
+In Figure 15, we present a comparative analysis of the X-ray luminosity light curves of GRB 230204B with a sample of other GRBs that have measured redshifts (2024arXiv240904871G).
 The X-ray flux data for these GRBs were sourced from the Swift XRT light curve repository444https://www.swift.ac.uk.
@@
 Furthermore, the afterglow data collected across radio, optical, and X-ray wavelengths showed that GRB 230204B’s luminosity at various stages aligns with typical long-duration GRBs.
-Optical follow-up using multiple ground-based observatories revealed a rapid, early afterglow decay rate, which was more luminous than most GRBs at a similar epoch.
-Its high optical luminosity and bright afterglow suggest efficient energy transfer from the burst to the surrounding medium, or an environment conducive to high luminosity, possibly due to a structured jet.
+Optical follow-up using multiple ground-based observatories (mainly using the MASTER and BOOTES robotic telescopes) revealed a rapid, early afterglow decay rate, which was more luminous than most GRBs at a similar epoch.
+The contributions of the MASTER and BOOTES robotic telescope networks were indispensable for the early detection and characterization of the optical afterglow of GRB 230204B.
+The scientific impact of the early robotic observations extends beyond simple discovery and can be understood through these key aspects:
+The continuous monitoring by the MASTER and BOOTES networks over the first few hours post-burst revealed a smooth power-law decay without evidence of optical flare, reverse shock emission, or other early-time features that have been observed in other bright GRBs (e.g., GRB 080319B 2008Natur.455..183R, GRB 990123, 1999Natur.398..400A).
+This forward-shock-dominated evolution, established through early observations, allowing robust constraints on the microphysical parameters, including the circumburst density, jet viewing angle, jet opening angle, and magnetic field properties through broadband modeling (see Section 3.4).
+These parameters are fundamental to understanding the energetics and collimation of this exceptionally luminous GRB, underscoring the importance of robotic networks in the study of transient astronomical phenomena.
+In particular, without these rapid robotic measurements, the steep optical decay and its placement on luminosity–decay correlations would remain unconstrained.
+The high optical luminosity and bright afterglow of GRB 230204B indicate efficient energy transfer from the burst to the surrounding medium or an environment conducive to high luminosity.
+Our afterglow modeling suggests that this can be explained within the framework of a top-hat jet.
 Broadband afterglow modeling was conducted to investigate the GRB’s jet structure and surrounding environment.
@@
 Further studies of GRBs with similar spectral evolution and energy efficiency will help clarify the diversity of radiation mechanisms in GRBs and their implications for jet physics and energy dissipation.
-Overall, our analysis of GRB 230204B underscores its importance as a reference for studying the emission mechanisms, jet structures, and afterglow behaviors of long-duration GRBs.
-The findings contribute to the growing understanding of GRB diversity and emphasize the role of broadband follow-up observations in exploring the complex evolution of these powerful cosmic events.
-Future studies of similar events, using well-coordinated observations across the spectrum, will be crucial in further refining our understanding of GRB physics and jet composition.
+Overall, our analysis of GRB 230204B highlights its significance as a benchmark for studying the emission mechanisms, and afterglow evolution of long-duration GRBs.
+The results not only enrich the broader understanding of GRB diversity but also underscore the indispensable role of broadband, multi-wavelength follow-up observations in probing the complex physical processes governing these events.
+In particular, this case study demonstrates that the value of robotic telescope networks extends far beyond rapid discovery: their early-time, high-cadence data provide essential constraints on jet dynamics, circumburst environments, and emission mechanisms that would otherwise remain ambiguous.
+As next-generation and expanded robotic networks become operational, their systematic ability to capture the earliest phases of GRB afterglows will be crucial for refining jet models, improving physical parameter estimates, and advancing our overall understanding of GRB physics and relativistic jet composition.
 Appendix A Figures
+| Model | Top hat | Gaussian | Power-law |
+| 0.008 | 0.016 | 0.006 | |
+| 54.322 | 54.087 | 54.620 | |
+| 2.125 | 2.070 | 2.249 | |
+| -0.315 | -0.069 | -0.221 | |
+| -2.168 | -3.066 | -3.265 | |
+| -1.376 | -0.684 | -1.033 | |
+| 1.0 | 1.0 | 1.0 |
 | Tstart | Tstop | Flux | DICBand | Flux | DICCPL | DIC | |||||
@@
 | 4028.6 | Clear | 14.45 | 0.08 | BOOTES | This work |
+| 27405.6 | Clear | 19.6 | – | BOOTES | This work |
+| 36280.3 | 17.7 | – | BOOTES | This work | |
+| 37487.6 | Clear | 19.7 | – | BOOTES | This work |
+| 37558.0 | 19.4 | – | BOOTES | This work | |
+| 72995.8 | 19.3 | – | BOOTES | This work | |
+| 74544.8 | 19.4 | – | BOOTES | This work | |
 | 101119.0 | 21.09 | 0.09 | DOT | This work | |
@@
 | 200738.2 | 21.67 | 0.12 | DOT | This work | |
-| 105611.8 | 22.27 | – | DOT | This work | |
+| 105611.8 | 21.4 | – | DOT | This work | |
 | 95569.7 | 18.94 | UVOT | This work | ||